### v1

# A Survey on Large Language Models with Multilingualism: Recent Advances and New Frontiers

Abstract
The rapid development of Large Language Models (LLMs) demonstrates remarkable multilingual capabilities in natural language processing, attracting global attention in both academia and industry. To mitigate potential discrimination and enhance the overall usability and accessibility for diverse language user groups, it is important for the development of language-fair technology. Despite the breakthroughs of LLMs, the investigation into the multilingual scenario remains insufficient, where a comprehensive survey to summarize recent approaches, developments, limitations, and potential solutions is desirable. To this end, we provide a survey with multiple perspectives on the utilization of LLMs in the multilingual scenario. We first rethink the transitions between previous and current research on pre-trained language models. Then we introduce several perspectives on the multilingualism of LLMs, including training and inference methods, model security, multi-domain with language culture, and usage of datasets. We also discuss the major challenges that arise in these aspects, along with possible solutions. Besides, we highlight future research directions that aim at further enhancing LLMs with multilingualism. The survey aims to help the research community address multilingual problems and provide a comprehensive understanding of the core concepts, key techniques, and latest developments in multilingual natural language processing based on LLMs.
11footnotetext: Equal Contribution. Contact e-mail: kyhuang@bjtu.edu.cn; fengran.mo@umontreal.ca
1 Introduction
With the rapid development of artificial intelligence (AI), the advent of large language models (LLMs) such as GPT-3.5 Ouyang et al. (2022), GPT-4 Achiam et al. (2023), and LLaMA Touvron et al. (2023) has emerged as groundbreaking technologies, revolutionizing the field of natural language processing (NLP). LLMs have pushed the boundaries of what was previously thought possible with a “Prompt” style Liu et al. (2023a). Their capability to understand and generate human-like text has achieved state-of-the-art performance in various downstream tasks such as machine translation Manakhimova et al. (2023); Hendy et al. (2023); Khatri et al. (2023), text summarization Luo et al. (2023a); Wang et al. (2023a), and sentiment analysis Sudirjo et al. (2023); Fatouros et al. (2023). Besides, the ability of LLMs to adapt and learn from vast amounts of data has made them indispensable tools for researchers, developers, and businesses across diverse industries Fan et al. (2023); Goyal et al. (2024). Importantly, as AI continues to evolve, the impact of LLMs on our society and technology is poised to grow even further, opening up new opportunities and challenges in the realm of natural language understanding and generation.
FIGURE 1: An illustration of the training process of LLMs with a fail case in each phase caused by multilingualism. Due to the long context of the shown case, we present only the key parts.
The existing LLMs are typically based on transformer architectures Vaswani et al. (2017) and trained on massive data that consists of a mixture of different languages. As shown in Figure 1, the training process of LLMs mainly consists of three stages: unsupervised pre-training, supervised fine-tuning with instructions, and aligning algorithm via reinforcement learning from human feedback (RLHF) Ouyang et al. (2022). Although LLMs have achieved remarkable advancements, their application in multilingual scenarios is still limited, especially in extremely low-resource languages Lai et al. (2023). The reason is that the language distribution of the training data for LLMs is highly imbalanced and the quality varies across languages Ding et al. (2024). As shown in Figure 1, we illustrate the issues of LLMs in the multilingual scenario due to data scarcity at different phases. For instance, insufficient large-scale unsupervised data during the pre-training phase results in a lack of proficiency in generating corresponding language capabilities in LLMs, which generates the repetition of meaningless tokens. Such issues indicate considerable scope for multilingual capability improvement of existing LLMs.
To alleviate the issue of multilingualism, a common practice is to enhance the multilingual capabilities of the LLMs by incorporating corresponding multilingual data at each training stage Yang et al. (2023a); Le Scao et al. (2022); Wang et al. (2023b). However, the LLMs in the literature are mostly large-scale data-driven and face the following language issues in the multilingual scenario: (1) Knowledge Transfer: Several studies Zhang et al. (2023a); Wang et al. (2023c); Zhang et al. (2024a) have demonstrated the necessity of employing appropriate data to unleash the multilingual potential of LLMs. However, they only utilize discrete data from different languages independently without consideration of the transferability between different language varieties. Thus, the existing LLMs cannot perform well on low-resource languages Chang et al. (2023); Li et al. (2024a) and limit the number of supported languages. (2) Knowledge Accumulation: The “data island”***The data island means that data has non-existent or limited external connectivity in different storages. https://en.wikipedia.org/wiki/Data_island exists in practical scenarios due to the limitation of resource availability. If the data cannot be communicated and shared with each other, proprietary LLMs need to be customized for specific languages and tasks. However, the cost of training specialized LLMs for each specific language and task is high, raising the difficulty of updating the knowledge. Besides, the general knowledge inherited in LLMs might also be forgotten during continual training, leading to catastrophic forgetting Kirkpatrick et al. (2017). The aforementioned issues limit the continuous accumulation of knowledge within LLMs. (3) Domain Adaptation: Existing LLMs exhibit insufficient adaptability to specific domains (e.g., medical and finance) in multilingual scenarios. The domain-customized models like BioGPT Luo et al. (2022) and FinBERT Liu et al. (2021) based on domain-specific corpora are mostly English-centric. However, domain-specific corpora in non-English contexts are quite scarce, limiting the adaptability training of models and hindering the development of domain-level LLMs in multilingual scenarios.
As LLMs are deeply integrated into various applications that need context comprehension and generation in multilingual scenarios, the requirement for LLMs capable of effectively and efficiently understanding and generating sequences in a language-agnostic manner becomes increasingly indispensable and urgent. Consequently, researchers have devoted significant efforts to enhancing the practicality of LLMs in multilingual scenarios from various perspectives, including training procedure (Section 3), optimization on the inference (Section 4), security in multilingual situations (Section 5) and multi-domain (Section 6). Furthermore, due to the paradigm shift from “Pre-train, Fine-tune” to “Pre-train, Prompt, Predict” in multilingual scenarios (Section 2), there has been a subtle change in the definition of multilingualism, leading to challenges in the development of the multilingual community. A series of multilingual approaches to facilitate LLMs with varying focuses have merged while lacking systematic categorization and comparison, raising the challenge of developing practical applications for specific language needs. Thus, new systematic literature and standardized definitions for introducing and comparing existing LLMs from a multilingual perspective are desirable.
As the field of LLMs is developing rapidly within the AI research community, some recent surveys try to summarize these developments to provide future guidance. Xu et al. Xu et al. (2024a) propose the first survey on multilingual LLMs from three aspects, which primarily analyzes the issues of data misalignment and bias, lacking exploration from the paradigm perspective. Different from them, we conduct a structured taxonomy and comprehensive review from several perspectives. A more recent survey Qin et al. (2024) closely resembles our study, but it only classifies in terms of alignment, which is less comprehensive than ours. In particular, we not only meticulously examine the multilingual capabilities and training methods of current LLMs, but also thoroughly investigate how to uncover the potential of LLMs. In this survey, we introduce the concept of “multilingual LLMs”, and provide a comprehensive and systematic survey of existing LLMs that have remarkable multilingual capabilities. We offer categorization, comparative analysis, and multi-perspective exploration for these models, evaluating their applicability and limitations, and providing practical recommendations for their effective real-world utilization. Additionally, we discuss some useful datasets and benchmarks related to multilingualism. We also present recommendations for future research. The main contributions are as follows:
  • A structured taxonomy. We rethink the transitions between previous and current research on LLMs, providing a systematic comparison and standardized definitions for multilingual LLMs. A broad overview of the field is presented with a structured taxonomy that categorizes existing studies (Figure 3).
  • Comprehensive reviews. We present a comprehensive investigation from several perspectives for the multilingualism of LLMs, including training and inference methods, model security, multi-domain with language culture, and usage of datasets.
  • Future directions. We identify the key challenges and provide potential solutions to advance the frontier for each summarized research direction, which is useful in enhancing the multilingual capabilities of LLMs,
  • A growing repository. Considering the rapid growth of the research of LLMs, we have established a repository to gather relevant literature in this specific multilingual domain and will continuously update it to maintain the latest advancementshttps://github.com/kaiyuhwang/MLLM-Survey.
Survey organization. The rest of this survey is organized as follows: In Section 2, we review the transition from pre-trained models to large language generative models in the multilingual scenario. In Section 3, we organize relevant models with various architectures, datasets, and training paradigms. Moreover, in Section 4, we investigate various multilingual inference strategies to harness the potential of LLMs for better accomplishment of multilingual tasks. In Section 5, we explore the security of LLMs with multilingual strategies, which is considered to be crucial for LLMs. In Section 6, we discuss the multi-domain issue in the multilingual real-world scenario. In addition, we present several available datasets for multilingual LLMs in Section 7 and highlight the benchmark and evaluation in Section 8. We discuss the cross-lingual bias evaluation and elimination in Section 9. Finally, we conclude this survey in Section 10.
2 Preliminary
In this chapter, we first give a general definition of multilingual models, then describe the background of pre-trained language models (PLMs), and finally introduce the paradigm shift from “Pre-train, Fine-tune” to “Pre-train, Prompt, Predict” in the age of pre-trained language models.
FIGURE 2: An overview of representative LLMs and mPLMs in recent years. The illustration consists of one tree that shows the transition of two paradigms (“Pre-train, Fine-tune”$\rightarrow$“Pre-train, Prompt, Predict”), including three model’s architectures (encoder-only, decoder-only, and encoder-decoder) and four new frontiers for LLMs with multilingualism.
2.1 Multilingual Models
The longstanding goal of multilingual models is to develop a universal model, capable of providing high-quality performance on any languages and tasks Huang et al. (2023a). Generally, multilingual models are trained to maximize the optimization of a mix of examples drawn from multiple language corpora Johnson et al. (2017). The benefits of multilingual models are mainly twofold. First, multilingual models can handle several languages in a single model, facilitating the knowledge transfer from high-resource languages to related but low-resource languages Huang et al. (2023b), where even languages that have never been trained before Zhang et al. (2020). Second, a single model can support multiple languages instead of training multiple language-specific models Lin et al. (2021a); Escolano et al. (2021), reducing the cost of maintaining models.
2.2 Pre-Trained Language Models
PLMs are deep neural networks initially trained on extensive unlabeled corpora, then could be adaptable to specific tasks. Existing research demonstrates that PLMs grasp and store substantial linguistic knowledge within their parameters Min et al. (2023); Li et al. (2020); Hu et al. (2023). Consequently, leveraging PLMs holds promise for enriching language comprehension and enhancing multilingual performance.
Most of the PLMs adopt the remarkable Transformer architecture Vaswani et al. (2017) as their backbone. The existing PLMs can be categorized into three typical architectures, constructed upon the Transformer encoder (e.g., BERT Devlin et al. (2018)), decoder (e.g., GPT Radford et al. (2018)), and encoder-decoder (e.g., BART Lewis et al. (2020a)), respectively. With the corresponding architecture, PLMs are usually improved based on training data sizes, parameter sizes and training strategies Li et al. (2022). For instance, GPT-3 Brown et al. (2020) is developed based on the GPT-2 Radford et al. (2019) by expanding the model size (from 1.5B to 175B) and the scale of training data, which improves performance across various language tasks. Recent research also indicates that scaling up model parameters enhances performance of PLMs Kaplan et al. (2020), leading to the emergence of large-scale PLMs such as GPT-3 (175B) Brown et al. (2020), OPT (175B) Zhang et al. (2022a), PaLM (540B) Chowdhery et al. (2023) and Switch-Transformers (1.6T) Fedus et al. (2022), boasting billions or even trillions of parameters. Besides, a new training strategy “translation language model (TLM)” is designed specifically for enhancing multilingual capabilities tasks compared to the masked language model (e.g., BERT Devlin et al. (2018)), involving translation between multiple languages Conneau and Lample (2019).
2.3 Multilingual Paradigm Transition
As shown in Figure 2, multilingual pre-trained language models (PLMs) are dominated by small language models (i.e., parameters less than 7B). They are designed for specific tasks, such as multilingual neural machine translation (NMT) Johnson et al. (2017); Aharoni et al. (2019); Firat et al. (2017), multilingual question answering (QA) Loginova et al. (2018, 2021); Ruder and Sil (2021) and multilingual reasoning Lin et al. (2021b); Fang et al. (2022); Bang et al. (2023), etc. To indicate the language of target tasks, a prepending language token is appended to each source sentence Johnson et al. (2017); Conneau and Lample (2019). Take the multilingual NMT for example, formally, given the source sentence $\mathbf{x^{\prime}}=(x_{1},x_{2},...,x_{n})$, where each $x_{i}$ is a token in the source language, the modified source sentence is represented as $\mathbf{x}=(l_{i},x_{1},x_{2},...,x_{n})$. Here, $l_{i}$ denotes the target language. Correspondingly, the target sentence is represented as $\mathbf{y}=(y_{1},y_{2},...,y_{m})$, where each $y_{j}$ represents a token in the target language. The probability of a target sentence is given by:
$p(\mathbf{y}|\mathbf{x};\theta)=\prod^{J}_{j=1}p(y_{i}|\mathbf{\mathbf{y}_{<j},\mathbf{x};\theta})$ (1)
where $\theta$ is a set of trainable parameters, $\mathbf{y}_{<j}$ are the generated words before the $j$-th step. This approach that prepends language identification is suitable for specific tasks and language situations.
However, the current model paradigm has gradually shifted to Artificial General Intelligence (AGI) which possesses the ability like human intelligence to perform a wide range of tasks Ouyang et al. (2022); Achiam et al. (2023); Yang et al. (2023b), including understanding, learning, cognition, communication and others. Unlike small language models, which are designed for specific tasks, AGI would have the capacity for generalization and adaptation to novel situations, potentially surpassing human capabilities in various domains. With the increasing scale of language models, LLMs improve the performance of downstream tasks compared to small language models Kaplan et al. (2020), which is closer to the goal of AGI. For the sake of generality, the task form becomes general tasks instead of specific tasks, and prepending the language is not allowed. Hence, most existing LLMs typically do not specify whether they function as multilingual models or detail the number of languages they support. We consider that if a certain amount of multilingual data is used in the training corpus, it can be regarded as a multilingual LLM. Among the existing models, only a few are explicitly called “multilingual LLM” (e.g., BLOOM Le Scao et al. (2022), BayLing Zhang et al. (2023a), etc), yet there are also other models that possess multilingual capabilities Jiang et al. (2023); Wang et al. (2023c). To distinguish LLMs from multilingual LLMs, we propose a definition of “LLMs with multilingualism”.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, calign=edge midpoint, , [LLMs with Multilingualism, text width=1.7cm, fill=black!10 [Representative Models §3, text width=1.8cm, for tree=fill=red!20 [Training from scratch §3.1, text width=1.8cm [ GPT-3 Brown et al. (2020); mT5 Xue et al. (2020); ByT5 Xue et al. (2022); Gopher Rae et al. (2021); LaMDA Thoppilan et al. (2022); OPT Zhang et al. (2022a); PaLM Chowdhery et al. (2023); mGPT Shliazhko et al. (2022); GPT-3.5 Ouyang et al. (2022); XGLM Lin et al. (2022); LLaMA Touvron et al. (2023); GPT-4 Achiam et al. (2023); PANGU-$\sum$ Ren et al. (2023); Pythia Biderman et al. (2023); PaLM-2 Anil et al. (2023); InternLM Team (2023); PolyLM Wei et al. (2023a); LLaMA-2 Touvron et al. (2023); Baichuan-2 Yang et al. (2023c); Qwen Bai et al. (2023); Mistral Jiang et al. (2023); Gemini Team et al. (2023); TigerBot Chen et al. (2023a); YAYI-2 Luo et al. (2023b); DeepSeek Bi et al. (2024); Orion Chen et al. (2024a); TeleChat Abdin et al. (2024); Claude3 Anthropic (2024); InternLM2 Cai et al. (2024); Phi-3 Abdin et al. (2024); LLaMA-3 AI@Meta (2024), text width=7.0cm, node options=align=left ] ] [Continual Training §3.2, text width=1.8cm [ FLAN-T5 Chung et al. (2024); FLAN-PaLM Chung et al. (2024); ChatGLM Zeng et al. (2022); Alpaca Taori et al. (2023); ParroT Jiao et al. (2023); BigTrans Yang et al. (2023a); Vicuna Chiang et al. (2023); BayLing Zhang et al. (2023a); OpenChat Wang et al. (2023c); GLM-4 ZHIPU (2024), text width=7.0cm, node options=align=left ] ] ] [Inference Strategies §4, text width=1.8cm, for tree=fill=yellow!20 [Translation-Based, text width=1.8cm [ Intrator et al. Intrator et al. (2024); Liu et al. Liu et al. (2024a), text width=7.0cm, node options=align=left ] ] [Chain-of-Thought §4.3, text width=1.8cm [ Liu et al. Liu et al. (2024a); XLT Huang et al. (2023c); Shi et al. Shi et al. (2022a); Kim et al. Kim et al. (2023a); Suzgun et al. Suzgun et al. (2023); xCoT Chai et al. (2024), text width=7.0cm, node options=align=left ] ] [Retrieval Augmented Generation §4.4, text width=1.8cm [ Zhang et al. Zhang et al. (2023b); Shi et al. Shi et al. (2022b); Agrawal et al. Agrawal et al. (2023); Li et al. Li et al. (2023a); Li et al. Li et al. (2023b); Winata et al. Winata et al. (2023a); Garcia et al. Garcia et al. (2023); Ramos et al. Ramos et al. (2023); Kim et al. Kim et al. (2023b); Thakur et al. Thakur et al. (2024); Sennrich et al. Sennrich et al. (2023); Vernikos et al. Vernikos and Popescu-Belis (2024); Fu et al. Fu et al. (2024); Zeng et al. Zeng et al. (2024); He et al. He et al. (2024); Conia et al. Conia et al. (2023), text width=7.0cm, node options=align=left ] ] [Code-Switching §4.5, text width=1.8cm [ Zhang et al. Zhang et al. (2023c); Koto et al. Koto et al. (2024); Peng et al. Peng et al. (2023a), text width=7.0cm, node options=align=left ] ] ] [Security §5, text width=1.8cm, for tree=fill=blue!20 [Attack (Jailbreak) §5.1, text width=1.0cm, for tree=fill=blue!20 [Greedy Coordinate Gradient, text width=1.4cm [ Sitawarin et al. Sitawarin et al. (2024); Zou et al. Zou et al. (2023), text width=6.0cm, node options=align=left ] ] [Prompt-Based, text width=1.4cm [ Wei et al. Wei et al. (2024); Liu et al. Liu et al. (2023b); Shen et al. Shen et al. (2023); Deng et al. Deng et al. (2024); Li et al. Li et al. (2024b); Liu et al. Liu et al. (2023c); Jin et al. Jin et al. (2024), text width=6.0cm, node options=align=left ] ] [Multilingual, text width=1.4cm [ Shen et al. Shen et al. (2024); Deng et al. Deng et al. (2023); Puttaparthi et al. Puttaparthi et al. (2023); Yong et al. Yong et al. (2023); Xu et al. Xu et al. (2023); Li et al. Li et al. (2024c); Yuan et al. Yuan et al. (2024); Huang et al. Huang et al. (2023d), text width=6.0cm, node options=align=left ] ] ] [Defense §5.2, text width=1.0cm, for tree=fill=blue!20 [Open-Source Models, text width=1.4cm [ Robey et al. Robey et al. (2023); Deng et al. Deng et al. (2023); Li et al. Li et al. (2024c); Zhou et al. Zhou et al. (2024a); GUARD Jin et al. (2024), text width=6.0cm, node options=align=left ] ] [Closed-Source Models, text width=1.4cm [ Jain et al. Jain et al. (2023); Wu et al. Wu et al. (2024); Li et al. Li et al. (2023c), text width=6.0cm, node options=align=left ] ] ] ] [Multidomain §6, text width=1.8cm, for tree=fill=green!20 [Medical §6.1, text width=1.8cm [ KBioXLM Geng et al. (2023); BioMistral Labrak et al. (2024); MMedLM2 Qiu et al. (2024a); Apollo Wang et al. (2024a); L2M3 Gangavarapu (2024); Medical mT5 Garc´ıa-Ferrero et al. (2024), text width=7.0cm, node options=align=left ] ] [Legal §6.2, text width=1.8cm [ LEXTREME Niklaus et al. (2023a); Brugger et al. Brugger et al. (2023); Christen et al. Christen et al. (2023); Baumgartner et al. Baumgartner et al. (2024); Niklaus et al. Niklaus et al. (2023b); LegalLAMA Chalkidis et al. (2023); Trautmann et al. Trautmann et al. (2022), text width=7.0cm, node options=align=left ] ] ] [Datasets, text width=1.8cm, for tree=fill=pink!20 [Corpus §7, text width=1.8cm [ Amazon intent mteb (a); Amazon reviews Keung et al. (2020a) ; Bactrian-x Li et al. (2023d); Biblenlp eBible.org ; Bloom-lm Le Scao et al. (2022); CC100 Conneau et al. (2020a) Wenzek et al. (2020); CulturaX Nguyen et al. (2023); GPT-4 Prompts erfanzar ; Guanaco Dettmers et al. (2023); HPLT de Gibert et al. (2024) ; IWSLT 2017 Cettolo et al. (2017); mC4 Raffel et al. (2019); Mewsli-x Ruder et al. (2021a); Minds14 Gerz et al. (2021) ; Miracl Zhang et al. (2022b) ; MLDR Chen et al. (2024b); MMedC Qiu et al. (2024b); MQA De Bruyn et al. (2021); Multi-sentiments tyqiangz ; Multiconer Fetahu et al. (2023a) Fetahu et al. (2023b); Nomiracl Thakur et al. (2023) Thakur et al. (2024);Open Subtitles Lison and Tiedemann (2016) ; OSCAR Abadji et al. (2022); Para-pat Soares et al. (2020) ; Project Gutenberg Gutenberg ; ShareGPT RyokoAI ; SREDFM Huguet Cabot et al. (2023); TED Talks Qi et al. (2018); TED-talks-iwslt Cettolo et al. (2012); Toxi-text FredZhang ; UD Nivre et al. (2020); Wikiann Pan et al. (2017) Rahimi et al. (2019); Wikipedia Foundation ; Wit Base Srinivasan et al. (2021); xP3 Muennighoff et al. (2022a), text width=7.0cm, node options=align=left ] ] [Benchmark §8, text width=1.8cm [ Afrisent Muhammad et al. (2023); ASPEN Razumovskaia et al. (2024) ; BELEBELE Bandarkar et al. (2023); BioInstructQA BioMistral ; Bucc-bitext-mining mteb (b); Cross-Sum Bhattacharjee et al. (2023); Crossmodal-3600 Thapliyal et al. (2022); Exams Hardalov et al. (2020); Fairlex Chalkidis et al. (2022a); FLORES-200 Team (2022) Goyal et al. (2022) Guzmán et al. (2019); GEOMLAMA Yin et al. (2022); Humaneval-XL Peng et al. (2024); M-Hellaswag Dac Lai et al. (2023); M-MMLU Dac Lai et al. (2023); M3Exam Zhang et al. (2023d); M3LS Verma et al. (2023); MARC Keung et al. (2020b); MasakhaNER Adelani et al. (2021); Masakhanews Adelani et al. (2023a); MASSIVE FitzGerald et al. (2022); MaXM Changpinyo et al. (2023); MEGA Ahuja et al. (2023); MEGAVerse Ahuja et al. (2024); Mela Zhang et al. (2024b); MGSM Shi et al. (2022a); MLQA Lewis et al. (2020b); MMedBench Qiu et al. (2024c); Multi-CoNER Malmasi et al. (2022); MULTIEURLEXDOC Chalkidis et al. (2021); Multilingual-Fig-QA Kabra et al. (2023); NusaX Winata et al. (2023b); ODEX Wang et al. (2023d); OPUS-100 Zhang et al. (2020); Paws-X Yang et al. (2019); Pmindiasum Urlana et al. (2023); PRESTO Goel et al. (2023); SEAHORSE Clark et al. (2023); Sib200 Adelani et al. (2023b); SMiLER Seganti et al. (2021); STSB-multi-mt May (2021); Tatoeba-mt Tiedemann (2020); Tydip Srinivasan and Choi (2022); TyDiQA Clark et al. (2020); Universal Dependencies Nivre et al. (2020) ; X-CLAIM Mittal et al. (2023); X-RiSAWOZ Moradshahi et al. (2023); XCOPA Ponti et al. (2020); XCSQA Lin et al. (2021b); xDial-Eval Zhang et al. (2023e); XGLUE Liang et al. (2020); XL-SUM Hasan et al. (2021); XNLI Conneau et al. (2018); XQuAD Artetxe et al. (2020); XSEMPLR Zhang et al. (2023f); XStoryCloze Lin et al. (2022); XTREME Hu et al. (2020); XTREME-R Ruder et al. (2021b); XWinograd Tikhonov and Ryabinin (2021); XCSR Lin et al. (2021c), text width=7.0cm, node options=align=left ] ] ] ]
FIGURE 3: A structured taxonomy of LLMs with multilingualism which categorizes current studies.
3 Large Language Models with Multilingual Capability
Name Release Time Params Affiliation Base Available Support Languages
GPT-3 Brown et al. (2020) 20-05 13B, 175B OpenAI GPT-2 Closed -
mT5 Xue et al. (2020) 20-10 13B Google T5 Open 101
ByT5 Xue et al. (2022) 21-05 13B Google T5 Open 101
Gopher Rae et al. (2021) 21-12 280B DeepMind - Open -
LaMDA Thoppilan et al. (2022) 22-01 137B Google - Open -
OPT Zhang et al. (2022a) 22-04 175B Meta - Open -
PaLM Chowdhery et al. (2023) 22-04 8B, 62B, 540B Google - Closed 124
mGPT Shliazhko et al. (2022) 22-04 13B - GPT-3 Open 61
BLOOM Le Scao et al. (2022) 22-07 176B BigScience - Open 46
FLAN-T5 Chung et al. (2024) 22-10 11B Google T5 Open 60
FLAN-PaLM Chung et al. (2024) 22-10 8B, 62B, 540B Google PaLM Open 60
ChatGLM Zeng et al. (2022) 22-10 130B ZHIPU GLM Open 2
GPT-3.5 Ouyang et al. (2022) 22-11 - OpenAI GPT Closed -
XGLM Lin et al. (2022) 22-11 7.5B Meta AI - Open 30
LLaMA Touvron et al. (2023) 23-02 7B, 13B, 33B, 65B Meta - Open -
GPT-4 Achiam et al. (2023) 23-03 - OpenAI - Closed -
Alpaca Taori et al. (2023) 23-03 7B StandFord LLaMA Open -
PANGU-$\sum$ Ren et al. (2023) 23-03 1085B Huawei - Closed 26
Pythia Biderman et al. (2023) 23-04 12B EleutherAI - Open -
ParroT Jiao et al. (2023) 23-04 7B Tencent Alpaca Open -
BigTrans Yang et al. (2023a) 23-05 13B - LLaMA Open 102
PaLM-2 Anil et al. (2023) 23-05 340B Google PaLM Closed -
Vicuna Chiang et al. (2023) 23-06 13B LMSYS LLaMA Open -
InternLM Team (2023) 23-06 104B Shanghai AI Laboratory - Open -
BayLing Zhang et al. (2023a) 23-06 7B, 13B ICT/CAS - Open -
PolyLM Wei et al. (2023a) 23-07 13B DAMO Academy - Open 18
LLaMA-2 Touvron et al. (2023) 23-07 7B, 13B, 34B, 70B Meta LLaMA Open -
Baichuan-2 Yang et al. (2023c) 23-09 7B, 13B Baichuan - Open -
Qwen Bai et al. (2023) 23-09 7B, 14B Alibaba - Open -
OpenChat Wang et al. (2023c) 23-09 7B, 13B Tsinghua LLaMA Open -
Mistral Jiang et al. (2023) 23-10 7B Mistral - Open -
Gemini Team et al. (2023) 23-12 - DeepMind - Closed 40+
TigerBot Chen et al. (2023a) 23-12 7B, 13B, 70B, 180B Tiger - Open -
YAYI-2 Luo et al. (2023b) 23-12 30B IACAS - Open -
DeepSeek Bi et al. (2024) 24-01 67B DeepSeek AI - Open -
GLM-4 ZHIPU (2024) 24-01 - ZHIPU GLM Close -
Orion Chen et al. (2024a) 24-01 14B OrionStar - Open 8+
TeleChat Abdin et al. (2024) 24-01 7B, 12B China Telecom - Open -
Claude3 Anthropic (2024) 24-03 - Anthropic - Closed -
InternLM-2 Cai et al. (2024) 24-03 7B, 20B Shanghai AI Lab Open -
LLaMA-3 AI@Meta (2024) 24-04 8B, 70B Meta - Open 30+
Phi-3 Abdin et al. (2024) 24-04 7B, 14B Microsoft - Open -
Table 1: An overview of representative LLMs (trainable parameters greater than 7B) that have certain multilingual capabilities in recent three years, including their release time, parameters, affiliation, base model, availability, and support languages.
Based on the training paradigm, this survey divides existing multilingual LLMs into two categories: (1) the foundational LLMs trained from scratch and (2) the continually trained LLMs on top of the foundational models. In Table 1, we present the statistics of the representative LLMs with certain multilingual capabilities in the past three years.
3.1 Training from Scratch
To obtain a language model with multilingual capability, a common practice is to leverage all available data in different languages for training. The language sampling algorithms are usually applied to control the importance of each language. As shown in Figure 2, multilingual pre-trained language PLMs before 2022 mainly consisted of two structures: (1) encoder-only designed for natural language understanding tasks and (2) encoder-decoder designed for natural language generation tasks. The parameters of these PLMs are not comparable with the existing LLMs. Recent studies show that according to the scaling law Kaplan et al. (2020); Muennighoff et al. (2024); Aghajanyan et al. (2023), the scale of the parameter has significant impacts on the performance of the models, i.e., the larger models lead to better performance. The researchers also observe that when the parameter scale of the language models exceeds a certain level, they not only achieve significant performance improvements but also find some special capabilities not observed within small-scale PLMs Kaplan et al. (2020); Muennighoff et al. (2024). To distinguish the difference in parameter scale, the research community specifies the term “LLMs” compared with PLMs based on a significant scale Zhao et al. (2023).
[…]
In addition to constructing multilingual data in the pre-training stage, the cross-lingual capabilities can also be improved by some specific methods (e.g., XLM Conneau and Lample (2019), XLM-R Conneau et al. (2019), XLM-V Liang et al. (2023), XGLM Lin et al. (2022), PaLM-2 Anil et al. (2023), PolyLM Wei et al. (2023a), mLongT5 Uthus et al. (2023)). For instance, the series of XLM Conneau and Lample (2019); Conneau et al. (2019); Liang et al. (2023) proposes a translation language modeling (TLM) pre-training task, utilizing parallel multilingual data to improve cross-lingual model pre-training, and it extends MLM by using batches of parallel sentences instead of consecutive sentences. Moreover, the sampling strategies are important to balance the proportion of multiple languages in training data, which facilitates models to learn multilingual representation Lin et al. (2022); Anil et al. (2023); Chung et al. (2023). PoLyLM Wei et al. (2023a) integrates bilingual alignment within the training corpus and implements a curriculum learning strategy to maintain a balanced representation of various languages during the pre-training phase. The model proposes a multilingual self-instruct approach, enabling the automatic generation of diverse multilingual instructions for fine-tuning the model.
An inevitable part of the multilingual learning paradigm for PLMs is to deal with vocabulary Dabre et al. (2020); Garcia et al. (2021); Huang et al. (2022). The “out-of-vocabulary (OOV)” tokens hurts the translation performance naturally for multilingual learning Zhang et al. (2022c). To alleviate this issue, ByT5 Xue et al. (2022) is a token-free model that operates directly on raw text (bytes or characters) which simplifies the process of different languages.
3.2 Continual Training
Another way to improve the multilingual capabilities of LLMs is continual training, which involves updating the model with new data rather than training a model from scratch. The main idea is to transfer knowledge from the foundation model and inject additional multilingual capability via the updated data, which does not require excessive computational or data resources and thus reduces the training cost.
This kind of training method utilizes existing monolingual or multilingual models for enhancing the specific languages ability (e.g., Chinese-LLaMA Cui et al. (2023a), Chinese-Alpaca Cui et al. (2023a), Chinese-Mixtral HIT-SCIR (2024), CPM-2 Zhang et al. (2021), LLaMAntino Basile et al. (2023), FinGPT Luukkonen et al. (2023), Sabiá Pires et al. (2023), Bode Garcia et al. (2024)) or extending to a wider range of languages, such as BigTrans Yang et al. (2023a), Glot500 ImaniGooghari et al. (2023) and others Ebrahimi and Kann (2021); Alabi et al. (2022); Muennighoff et al. (2022b). A branch of work is to directly continually train the foundation models without any other techniques Luukkonen et al. (2023); Yang et al. (2023a); Alabi et al. (2022). For instance, Muennighoff et al. Muennighoff et al. (2022b) introduce xP3, a corpus consisting of tasks in 46 languages, and utilizes a cross-lingual multitask fine-tuning approach to fine-tune the BLOOM and mT5 on the newly created corpora, along with the English-only P3 corpus, resulting in the production of BLOOMZ and mT0Z. Considering the importance of vocabulary for multilingual models, prior work employs vocabulary extension to improve performance for low-resource languages ImaniGooghari et al. (2023). Another branch of work is to introduce external parameters and strategies to learn new languages for multilingual models Ebrahimi and Kann (2021). For instance, Chinese-LLaMA Cui et al. (2023a) integrates an additional 20K Chinese tokens into the original vocabulary to improve the Chinese ability, then employs the Low-Rank Adaptation (LoRA) approach to facilitate efficient training on Chinese corpus. CPM-2 Zhang et al. (2021) divides the pre-training into three phases, including the monolingual pre-training, bilingual pre-training, and mixture-of-expert pre-training. Multi-phase continual training, incorporating knowledge inheritance, offers a substantial reduction in computation costs compared to training from scratch. Moreover, Wang et al. Wang et al. (2022a) systematically investigate strategies aimed at reducing dependence on traditional language resources by leveraging bilingual lexicons to fine-tune pre-trained multilingual models for underrepresented languages.
3.3 Limitations and Future Directions on Training Paradigm
Although the existing LLMs demonstrate a certain superiority in non-English language, either by training from scratch or continuing training on top of a foundation model with extended language data, these approaches based on text knowledge only still face challenges in addressing more complex scenarios as listed below.
Low-Resource scenarios. The current training paradigm relies on large-scale data for the target language and task, e.g., supervised fine-tuning based on annotated data or continual pre-training based on unsupervised data. Both approaches require a substantial amount of data collection and thus constraints the effectiveness under low-resources or multi-scenario demands.
Knowledge conflict. As the foundation model expands the supported languages, its multilingual knowledge accumulates into the model through continuous new data available. The newly acquired knowledge from the new data would interfere with old knowledge stored in the parameters, and thus lead to a rapid decline in model performance. This is known as catastrophic forgetting Kirkpatrick et al. (2017), which might result in even completely forgetting previously learned knowledge.
Knowledge type. The existing multilingual NLP datasets, such as XNLI Conneau et al. (2018) and XQuAD Artetxe et al. (2020), are constructed based on the translated texts from the original English version datasets Rajpurkar et al. (2016); Yin et al. (2021). This is because of the high cost of obtaining multilingual data, especially in languages with limited annotations/translators. As a result, the essence of existing datasets lies in translation-based aligned knowledge, lacking cultural and domain knowledge in non-English contexts. This type of knowledge learned by the models reduces the distribution shift ability in the real-world scenario, leading to sub-optimal performance in specialized fields (e.g., finance and legislation).
To address the aforementioned problems and aim to improve the multilingual capabilities of existing LLMs, this survey proposes the following points for future exploration:
  • Training Strategies. LLMs are data-driven and follow the scaling law which suggests that increasing the model size can guarantee better generalization, higher accuracy, and improved ability to capture complex patterns from data. However, such a paradigm results in the existing training strategies primarily focusing on expanding diverse data, while overlooking other multilingual training techniques, such as sampling algorithms Arivazhagan et al. (2019); Wang et al. (2020); Wu et al. (2021), training objective optimization Yang et al. (2021); Pan et al. (2021), representation space Oncevay et al. (2020); Cheng et al. (2022); Stap et al. (2023), and others Zhang et al. (2020); Wang and Zhang (2022). The potential of these traditional multilingual training strategies has not been further explored in the current data effort paradigm in existing multilingual literature.
  • Architecture. Existing LLMs mainly rely on monolingual backbone models, such as the standard Transformer, rather than undergoing architectural changes specifically targeting multilingual capabilities. For the other specific tasks, for example, LLMs designed for long texts introduce variations in positional encoding Beltagy et al. (2020); Su et al. (2024), and those designed for multi-modal tasks enhance alignment between different modalities Peng et al. (2023b); Liu et al. (2024b) have been explored in some extend. Some advanced architectures like Mamba Gu and Dao (2023) have also shown remarkable performance and may become the backbone architecture for the next generation of LLMs. Thus, for multilingual tasks, it is essential to tailor the architecture with specific variations rather than merely augmenting data on the standard Transformer.
  • Sustainability. When trying to utilize new data to enhance existing LLMs, a practical approach is to allow LLMs to continually learn from the updated data, akin to how the human brain operates, instead of creating new models Huang et al. (2022, 2023a, 2023b). Essentially, when new language data comes up, we aim to continually extend the language support of LLMs and improve their corresponding language capabilities while preserving the performance of those languages that the model has already demonstrated good performance, as forming a lifelong/incremental learning paradigm.
4 Multilingual Inference Strategies
This chapter investigates the development of robust multilingual inference strategies that are crucial for deploying language models across varied linguistic environments.
4.1 Direct Inference in Multilingual Models
Model Inference XCOPA XStoryCloze BELEBELE XLSum XQuAD TyDiQA-GP
Acc Acc Acc Rouge-L F1 F1
PaLM2-S En-Pivot 87.3 96.4 76.7 23.7 67.2 81.6
Direct 89.7 96.8 77.8 26.8 70.7 83.8
PaLM2-L En-Pivot 89.6 97.8 84.3 25.4 78.7 81.0
Direct 93.4 99.1 88.4 28.0 85.9 83.0
Table 2: The performance of LLMs using different inference strategies in the multilingual scenario. The best score is highlighted in bold. $\dagger$” indicates the results are quoted from Intrator et al. (2024).
With the advent of LLMs, efforts to enhance the diversity of training corpora lead to the inclusion of multiple languages alongside English. This approach endows LLMs with inherent multilingual capabilities, which enables these models to engage in direct multilingual inference. In other words, the models can process the input in their native language without requiring translation into a pivot language. This capability is valuable as it maintains the authenticity of the linguistic and cultural nuances present in the original text, preventing semantic distortion or information loss that might otherwise occur during translation.
Recently, significant advancements have been made in enhancing the multilingual capabilities of LLMs, greatly increasing the number of languages covered and the performance in non-English inference. Advanced models such as GPT-4 Achiam et al. (2023) and PaLM-2 Anil et al. (2023) demonstrate remarkable multilingual capabilities and support hundreds of languages. The support of multiple languages enables direct multilingual inference to become feasible and come up with a current focal point of LLM research. As shown in Table 2, we investigate the comparison between direct inference and the pre-translation method which translates the prompt into a high-resource language (e.g., English and Chinese) before inference. The direct inference achieves better performance compared to the pre-translation method based on both PaLM2-S and PaLM2-L. The result demonstrates that these two LLMs have multilingual capabilities without the need to use English as a pivot for other language tasks. Moreover, without the requirement of the translation step, the direct inference approach reduces computational overhead and simplifies the processing pipeline with higher efficiency. The observations also confirm the benefits of direct inference, including the preservation of linguistic authenticity, enhanced processing efficiency, and improved performance in low-resource languages.
4.2 Pre-Translation Inference
The direct inference may not work for all LLMs, depending on their multilingual capacities. The existing LLMs usually perform better on those high-resource languages than the low-resource ones, because of the imbalance ratio within the training data. To enhance the performance on low-resource languages, pre-translation inference standardizes the input with various languages by translating them into a pivot high-resource language (e.g., English or Chinese) before querying the LLMs, which is based on the proficiency of the pivot language within the LLMs Huang et al. (2023c); Qin et al. (2023); Etxaniz et al. (2023). However, the guarantee of this method relies on high-quality translation services available, which are not necessarily true for most languages. As a result, the translation errors would accumulate and the final output of LLMs would become wrong. Besides, translation on a pivot language obscures the cultural and linguistic nuances of the original text, which might also lead to inaccurate results. As an exploratory work, Liu et al. Liu et al. (2024a) explore the performance comparison between direct inference using native languages and inference after translating into English for multilingual tasks. Although translation can enhance the multilingual reasoning performance of English-centric LLMs, for high-resource languages and advanced LLMs, reasoning in native languages is more effective. Besides, pre-translating to English is a practical approach in terms of current LLMs as they are predominantly trained on English data from their pilot experimental results. However, with the development of direct multilingual inference, the intermediary step may not be necessary, which could allow more authentic interactions with LLMs under multilingual scenarios.
4.3 Multilingual CoT
The Chain of Thought (CoT) is an effective approach to enhance the performance of LLMs in complex reasoning Sap et al. (2020); Yu et al. (2023a); Liu et al. (2023d), which has been extensively explored in existing studies, mainly focusing on English Wang et al. (2022b); Lyu et al. (2023); Wei et al. (2022a). Liu et al. Liu et al. (2024a) propose several inference strategies to investigate the effectiveness of prompting the LLMs in the multilingual scenario. For instance, “Native-CoT” indicates that asking questions in the native language with the instruction “Let’s think step by step.” is translated into the native language. “En-CoT” indicates that asking questions in the native language but instructing reasoning in English with “Let’s think step by step in English.” and “XLT” Huang et al. (2023c) indicates that translating questions into English and solving them step-by-step. “Trans” indicates that using the translation systems to convert questions into English and then solving them step-by-step. The experimental results demonstrate that the overall performance of the CoT instruction in English (En-CoT) is better than the one with instruction in native languages (Native-CoT). The results demonstrate that the CoT is still effective while using English to form the instructions for non-English queries. On the other hand, multilingual CoT attempts to enhance the reasoning capabilities of LLMs across multiple languages Shi et al. (2022a). The multilingual CoT approach is especially beneficial for complex reasoning tasks deeply embedded in specific cultural contexts, enabling more natural and intuitive problem-solving Kim et al. (2023a); Suzgun et al. (2023). The common practice of multilingual CoT is to prompt the LLMs to establish a step-by-step reasoning process in the original language of queries, which can preserve linguistic and cultural nuances. In addition, the results corresponding to different sizes of LLaMA-2-Chat attend to examine the impact of scaling law with different multilingual inference strategies.
The understanding of the syntax and semantics of various languages, such as idiomatic expressions and culture-specific references is challenging for existing LLMs. A recent study Chai et al. (2024) introduces a cross-lingual CoT reasoning instruction fine-tuning framework, which randomly replaces language fragments with those from low-resource languages, and mixes the original and target languages within a single query. Besides, this study creates multilingual CoT instruction training data, which can be used to supervise fine-tuning LLMs to reduce the performance gap across different languages. However, developing datasets for training and evaluating multilingual CoT capabilities requires a robust representation of linguistic diversity and detailed language-specific nuances, which should be further addressed.
4.4 Multilingual Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) is a methodology that integrates text generation with external knowledge retrieval, dynamically enhancing the quality of model response and accuracy by accessing relevant information Gao et al. (2023a); He et al. (2023a). This approach enables the model to utilize up-to-date or specialized knowledge in text generation, thereby increasing its practicality and reliability. A main branch of the multilingual RAG adopts to retrieve knowledge from the open domain and applies it to the in-context (i.e., augmented prompt) Zhang et al. (2023b); Shi et al. (2022b); Agrawal et al. (2023); Li et al. (2023a, b); Winata et al. (2023a); Garcia et al. (2023); Ramos et al. (2023); Kim et al. (2023b). In particular, Thakur et al. Thakur et al. (2024) propose NoMIRACL, a dataset across 18 languages, to evaluate the hallucination of LLM when given a piece of text in external retrieved knowledge. It consists of two parts, the non-relevant and the relevant subset, corresponding to the question and the passage which are non-related and related, respectively. The model is expected to answer “I don’t know” in the non-relevant subset, otherwise it is determined as a hallucination. In addition, hallucinations and off-target issues occur when incorporating LLMs with low-resource machine translation Sennrich et al. (2023), where the RAG can mitigate these issues via improving the translation quality for low-resource directions Vernikos and Popescu-Belis (2024); Fu et al. (2024); Zeng et al. (2024); He et al. (2024). Overall, the RAG can effectively facilitate LLMs to generate more reliable responses, alleviating the issues of hallucination and factual error without fine-tuning. However, achieving substantial enhancements solely through the RAG method for low-resource languages, where LLMs struggle, poses a considerable challenge. Meanwhile, it is also challenging to build retrievers that can be used in low-resource languages Conia et al. (2023).
4.5 Code-Switching
Module Model SA Aguilar et al. (2020) LID Aguilar et al. (2020) MT Srivastava and Singh (2022) Summarization Mehnaz et al. (2021)
Es-En (F1) Hi-En (F1) Hi-En (BLEU) Hi-En (ROUGE-L)
0-shot XGLM-7.5B 68.52 0.27 1.43 5.92
LLaMA-7B 51.28 0.44 1.44 3.37
GPT-3.5-turbo 75.64 78.17 27.64 25.07
5-shot XGLM-7.5B 61.06 16.91 3.28 5.18
LLaMA-7B 58.77 15.55 5.14 6.01
GPT-3.5-turbo 76.21 80.19 28.90 26.77
SFT XGLM-7.5B 80.32 80.06 28.11 -
LLaMA-7B 77.21 55.32 16.58 -
Table 3: The overall results on code-switching benchmark for LLMs. We report the 0-shot, 5-shot, and SFT performance. SFT only adapts to open-resource LLMs. “SA”, “LID”, and “MT” denote the Sentiment Analysis, the Language Identification, and the Machine Translation, respectively.
Code-switching refers to the phenomenon where communicators switch between two or more languages during linguistic interactions, based on contextual needs. This phenomenon is common in bilingual or multilingual communities, especially in spoken communication Zhang et al. (2023g); Huang et al. (2024). For example, in conversations involving Chinese-English or Spanish-English direction, speakers may freely switch between languages depending on the fluency, precision of expression, or emotional needs of the dialogue Doğruöz et al. (2021); Winata et al. (2021); Aguilar et al. (2020). Solving code-switching texts is an important and challenging task as the language IDs are not specified before LLMs inference Zhang et al. (2023c); Koto et al. (2024).
As shown in Table 3, we investigate the performance of LLMs on code-switching tasks. The results demonstrate that open-source LLMs cannot deal with the code-switching problem that needs to process multiple languages simultaneously under a single query. Multiple languages increase the complexity of processing and understanding language for models and LLMs need to continue to be supervised fine-tuned to accommodate simultaneous processing in this scenario. In addition, to improve the ability to deal with code-switching texts, researchers develop new pre-training techniques to enhance the understanding capacity of this linguistic phenomenon in LLMs Peng et al. (2023a). For instance, Das et al. Das et al. (2023) modify the MLM Devlin et al. (2018) to enhance model reasoning performance in code-switching environments. By employing code-switching corpora, it masks tokens at the boundaries between two languages, thus forcing the model to learn semantic nuances at the points of language transition. Additionally, this work introduces modifications to the model structure by adding residual connections from intermediate layers to the final layer. An auxiliary loss function based on the representations of intermediate layers is also proposed, which enhances the cross-lingual ability to understand multiple languages in a single sentence. However, the cost of annotated code-switching data is expensive and it is challenging to investigate the non-parametric or semi-parametric method to adapt the code-switching problems.
4.6 Limitations and Future Directions on Inference Strategies
The multilingual inference strategies exhibit a diverse range of characteristics. Although existing methods contribute to the performance of LLMs to some extent in the multilingual scenario, they still have many limitations as follows.
Universal inference paradigm. Previous PLMs are task-specific and language-specific, indicating that it is essential to specify the language ID. For instance, we need to assign the language ID of input and output respectively when leveraging the M2M-100 for machine translation Fan et al. (2021). However, it is not required in the existing paradigm of LLMs, such as GPT-4. Due to the universality of LLMs, an aim is to keep them as task-agnostic and language-agnostic as feasible. Current LLMs employ a standardized approach for inference across all languages (either direct inference, pre-translation inference, CoT, or RAG). However, as shown in the experiments and analyses above, the inference strategies exhibit varied performances when faced with different LLMs, tasks, and languages. Therefore, a flexible and universally applicable paradigm is desirable to uniformly handle the environments of all languages (i.e., high-/medium-/low-resource languages).
Language-Specific characteristics. Existing multilingual inference strategies are primarily adapted from monolingual (e.g., English) strategies, lacking exploration of language-specific characteristics. For example, the prompt engineering for a low-resource language is derived from English instructions, but these prompts are more aligned with the habits of English speakers instead of native speakers. Therefore, the approach might pose challenges in harnessing the potential of native languages during inference with LLMs.
Emergence ability. One important advantage of LLMs is the emergence ability that is informally defined as “the ability that does not exist in small models but appears in large models” Wei et al. (2022b). However, this advantage is not embodied significantly in multilingual scenarios. Thus, the performance of LLMs is far behind that of the English scenario, and difficult to outperform small language models. It is not only due to limitations during the training phase but also potentially influenced by inappropriate strategies during the inference phase.
Model collaboration. Existing inference pipelines are designed for single LLMs. Due to data scarcity in low-resource languages, it is challenging to address tasks for all languages within a single model Magueresse et al. (2020); Ranathunga et al. (2023). Considering that the smaller models excel at specific tasks or handling specific languages, which do not require large-scale data to optimize Miceli-Barone et al. (2017); Dabre et al. (2019); Bapna and Firat (2019). It is beneficial to leverage the strengths of both large and small models and solve tasks for different resource languages through model collaboration. The model collaboration that effectively bridges the large and small models is a direction yet to be explored for enhancing inference performance in the multilingual scenario.
5 Security of Multilingual Large Language Models
With the wide deployment of LLMs in various applications, increasing security concerns have emerged. This chapter introduces the security aspects of LLMs in the multilingual scenario, specifically exploring attack methods and the existing research on defense mechanisms. Since there are no clear definitions to distinguish whether an LLM is a multilingual model or not, this survey not only focuses on security issues specific to different languages but also provides a perspective on common security issues. The investigated methods work equally across all languages and can be easily transferred to multilingual scenarios, providing inspiring thoughts on future research.
Model Prompt-Based Jailbreak
JailBroken Wei et al. (2024) GPTFUZZER Yu et al. (2023b) AutoDAN Liu et al. (2023c) DeepInception Li et al. (2023e) ICA Wei et al. (2023b) PAIR Chao et al. (2023) ReNeLLM Ding et al. (2023)
GPT-3.5-turbo 100% 35% 45% 66% 0% 19% 87%
GPT-4-0613 58% 0% 2% 35% 1% 20% 38%
LLaMA-2-7B-Chat 6% 31% 51% 8% 0% 27% 31%
LLaMA-2-13B-Chat 4% 41% 72% 0% 0% 13% 69%
Vicuna-7B-v1.5 100% 93% 100% 29% 51% 99% 77%
Vicuna-13B-v1.5 100% 94% 97% 17% 81% 95% 87%
ChatGLM-3 95% 85% 89% 33% 54% 96% 86%
Qwen-7B-Chat 100% 82% 99% 58% 36% 77% 70%
Intern-7B 100% 92% 98% 36% 23% 86% 67%
Mistral-7B 100% 99% 98% 40% 75% 95% 90%
GCG series Multilingual Jailbreak
GCG Zou et al. (2023) Multilingual Deng et al. (2023) Cipher Yuan et al. (2024) CodeChameleon Lv et al. (2024)
GPT-3.5-turbo 12% 12% 100% 35%
GPT-4-0613 0% 0% 58% 0%
LLaMA-2-7B-Chat 46% 46% 6% 31%
LLaMA-2-13B-Chat 46% 46% 4% 41%
Vicuna-7B-v1.5 94% 94% 100% 93%
Vicuna-13B-v1.5 94% 94% 100% 94%
ChatGLM-3 34% 34% 95% 85%
Qwen-7B-Chat 48% 48% 100% 82%
Intern-7B 10% 10% 100% 92%
Mistral-7B 82% 82% 100% 99%
Table 4: An overview of Greedy Coordinate Gradient, Prompt-Based and Multilingual attack methods jailbreaking for LLMs on the AdvBench Zhou et al. (2024b). The evaluation method is consistent with the EasyJailbreak Zhou et al. (2024b) framework, which uses GPT-4-turbo-1106 as the scoring model and the evaluation prompts from GPTFUZZER Yu et al. (2023b).
5.1 Attack Methods
To explore the security of LLMs, a red team attack on LLMs is required. A red-team attack is a cybersecurity exercise where a group of ethical hackers, known as the red team, simulate real-world cyberattacks on an organization’s systems, networks, or infrastructure. The goal of a red-team attack is to identify vulnerabilities, weaknesses, and potential security breaches that malicious actors could exploit Mirkovic et al. (2008). A common practice is the “jailbreak” attack which typically refers to the unauthorized access or modification of the underlying code or functionality of models. Essentially, it involves breaking out of the constraints or limitations imposed by the design or usage policies of LLMs. It includes techniques to bypass security measures or enable functionalities that are not intended or permitted by the developers.
According to the criteria in existing studies Wu et al. (2024); Zhou et al. (2024b), the jailbreak methods on LLMs can be divided into three types: Greedy Coordinate Gradient (GCG) jailbreak Sitawarin et al. (2024); Zou et al. (2023), prompt-based jailbreak Wei et al. (2024); Liu et al. (2023b); Shen et al. (2023); Deng et al. (2024); Li et al. (2024b); Puttaparthi et al. (2023); Liu et al. (2023c); Jin et al. (2024) and multilingual jailbreak Shen et al. (2024); Deng et al. (2023); Puttaparthi et al. (2023); Yong et al. (2023); Xu et al. (2023); Li et al. (2024c); Yuan et al. (2024); Huang et al. (2023d). In this section, we focus on these three jailbreak methods, especially the multilingual jailbreak. The first two methods involve generic attacks on LLMs, and the latter emphasizes jailbreaking through multiple languages. All these methods are aimed at bypassing the security measures of LLMs to generate malicious information. Since most of the jailbreaking methods are customized, the effectiveness of each method varies depending on the specific LLMs. To provide a comprehensive comparison, we investigate the performance of different jailbreaking methods across various LLMs based on a unified evaluation framework, as shown in Table 4.
5.1.1 Greedy Coordinate Gradient Jailbreak
The Greedy Coordinate Gradient  Zou et al. (2023) is an attack method based on greedy algorithms and gradient metric design hints. It first creates a seed prompt, and then iteratively replaces the tokens. The best prompt for attacking the model is determined by calculating the gradient. Sitawarin et al. Sitawarin et al. (2024) propose GCG++ to increase the attack success rate (ASR) of the original GCG method on LLaMA-2-7B. On the basis of the original GCG, they replace the cross entropy loss with multi-class loss to avoid vanishing gradients in the $Softmax$, which is effective in attacking LLM. Another discovery is that LLMs such as LLaMA-2-7B are format sensitive, having a strong prior for predicting a space token at the beginning of the model response. GCG++ adds a simple format-ware target string (i.e., “[ASSISTENT]:” before the input prompt to force the model to output harmful content. Through the two changes, GCG++ increases the GCG attack success rate on LLaMA-2-7B from 56% to 80%.
5.1.2 Prompt-Based Jailbreak
Characteristic Attack Method Types
GCG series LLM-Based Rule-Based
The LLM type for optimizing prompt Target LLM Agent LLM -
The status of Jailbreak prompt Dynamic Dynamic Static
The form of generating new prompts Auto Auto Manual
Table 5: An overview of comparison among GCG, Prompt-based and Rule-based jailbreak methods. Target LLM and Agent LLM refer to the attacked LLM and other LLMs that are different from target LLM, respectively.
The prompt-based methods aim to mislead LLM into generating harmful content based on specifically designed templates. Among them, LLM-based Liu et al. (2023c); Mehrotra et al. (2023); Sitawarin et al. (2024); Chao et al. (2023); Deng et al. (2024); Yu et al. (2023b); Ding et al. (2023); Bhardwaj and Poria (2023) and rule-based Wei et al. (2024); Li et al. (2023e); Liu et al. (2023b); Shen et al. (2023); Puttaparthi et al. (2023) are the two common used methods for designing attacking prompts. The differences between GCG, LLM-based, and rule-based are presented in Table 5 with various aspects.
Rule-based jailbreak methods are those works that design specific prompts to formulate rules for LLMs Wei et al. (2024); Li et al. (2023e), or directly collect prompts from various channels such as websites and forums Liu et al. (2023b); Shen et al. (2023); Puttaparthi et al. (2023). Rule-based jailbreak methods rely more on human participation because these rules are customized based on empirical experience in specific scenarios. Thus, the produced prompts are effective in a narrow area but lack versatility and are static, which can be easily defended.
LLM-based jailbreak methods leverage the instances of LLMs to optimize existing attack prompts or generate new jailbreak prompts. Different from the GCG series approaches, the principle of LLM-based jailbreak methods acts as using an agent LLM to attack a target LLM. Existing studies  Liu et al. (2023c); Mehrotra et al. (2023); Sitawarin et al. (2024); Chao et al. (2023) usually take the agent LLM as an optimizer, scoring module or evaluator to optimize prompts for attackers. Some other methods Deng et al. (2024); Yu et al. (2023b); Ding et al. (2023); Bhardwaj and Poria (2023) use the generation capabilities of LLM to perform multiple operations such as rewriting and shortening feasible attack prompts. They also wrap harmful prompts with scenarios to avoid safe alignment. These operations based on LLMs can enable the original attack prompt to evade the security alignment policy.
5.1.3 Multilingual Jailbreak
Due to the unequal security alignment of LLMs across different languages, with higher security in high-resource languages compared to low-resource languages, the multilingual jailbreak methods Shen et al. (2024); Deng et al. (2023); Yong et al. (2023); Xu et al. (2023); Li et al. (2024c) attend to cheat the LLMs via the alignment vulnerabilities between high-resource and low-resource language. Some other methods Yuan et al. (2024); Huang et al. (2023d); Lv et al. (2024); Wei et al. (2024) use special encoding or ciphers to disguise harmful prompts to achieve jailbreak, which can be considered as a type of special language.
The existing studies on the potential security risks of LLMs, primarily focused on the usage of English scenarios. Under the multilingual scenario, Deng et al. Deng et al. (2023) reveal the presence of multilingual jailbreak challenges within LLMs, who divide potential risks into two scenarios: (1) unintentional and (2) intentional. The unintentional scenario involves users inadvertently bypassing the safety alignment by querying LLMs with non-English prompts, while the intentional scenario refers to users intentionally combining malicious instructions with multilingual prompts. Compared to high-resource languages, low-resource languages are approximately three times more likely to encounter harmful content in GPT-3.5 and GPT-4, and thus the LLMs are much easier to attack via low-resource languages. This could also be the reason why the model attack approaches are designed with low-resource languages within these two scenarios.
For the unintentional scenario, Shen et al. Shen et al. (2024) indicate that the bottleneck in cross-lingual alignment lies within the training stage of LLMs. Thus, they study the effect of instruction tuning with RLHF and supervised fine-tuning (SFT) on the HH-RLHF dataset Bai et al. (2022). The results show that although training with high-resource languages can improve model alignment, the effect of training with low-resource languages is still negligible. For the intentional scenario, various prompts are designed to bypass the security defenses of the LLMs via high-resource languages. In particular, Yong et al. Yong et al. (2023) successfully circumvent the safeguard of GPT-4 by translating unsafe English prompts into low-resource languages. They propose the LRL-Combined Attacks approach to achieve 79% ASR on the AdvBench dataset Zou et al. (2023). Xu et al. Xu et al. (2023) study a new type of black box jailbreak attack, Cognitive Overload, which is specifically designed for the cognitive structure and processes of LLMs. With Google Cloud API, they translate original English harmful instructions from AdvBench and MasterKey Deng et al. (2024) into 52 other languages and propose multilingual cognitive overload Sweller (1988, 2011) to hinder in-context learning and inference process when the knowledge exceeds the limited capacity of LLMs Szulewski et al. (2021). By using Google Cloud API to translate English to other 52 languages and nllb-200-distilled-1.3B to translate other non-English responses back to English, they conducted a multilingual version of their attack method. With multilingual Cognitive Overload, their results show that LLMs are more vulnerable to non-English adversarial prompts. The observation indicates that the further a language is from English, the more effective the malicious prompt conveyed is in attacking LLMs. Moreover, Li et al. Li et al. (2024c) conduct an extensive empirical study on multilingual jailbreak attacks, which develops a novel semantic preservation algorithm to create multilingual jailbreak prompts. With these prompts, they reveal patterns in multilingual jailbreak attacks and implement fine-tuning mitigation methods for defending against cross-lingual jailbreak attacks.
Password jailbreak. From the perspective of writing scripts, special encoding or password can be considered as a language conversion, i.e., an English prompt can be converted into a password representation by encryption. Though the password representation can also be converted to English or other languages via decryption, the encrypted prompts cannot be understood directly by humans but by the LLMs, and thus could lead to secure alignment fails. For example, Huang et al. Huang et al. (2023d) propose the generation exploitation attack to manipulate variations of decoding methods (e.g., greedy decoding and sampling-based decoding), which can easily disrupt model alignment because existing alignment procedures are based on default decoding settings. Vulnerabilities are exposed when configurations of decoding change slightly. Yuan et al. Yuan et al. (2024) propose a novel framework CipherChat to examine the generalizability of safety alignment to non-natural languages, where the user can chat with LLMs through cipher prompts. Furthermore, Lv et al. Lv et al. (2024) introduce a hypothesis for the safety mechanism of LLMs: intent security recognition followed by response generation. Following this hypothesis, they propose CodeChameleon to transform queries into decryptable formats with custom Python functions. This approach enables the modification of the original query and achieves state-of-the-art ASR on 7 LLMs. Besides, Wei et al. Wei et al. (2024) propose Jailbroken, which obfuscates the queries using base64 to bypass the safety training of LLMs. They observe that LLMs can understand base64 but cannot defend against attacks under base64.
Other jailbreak. Apart from the aforementioned methods, Shayegani et al. Shayegani et al. (2023) combine an image targeted towards toxic embeddings with generic prompts to accomplish the jailbreak, which utilizes four embedding space target strategies to poison the vision encoder. Their attacks achieve 87% and 63.3% ASR on two vision-language models, LLaVA Liu et al. (2024b) and LLaMA-Adapter V2 Gao et al. (2023b), respectively. Rando et al. Rando and Tramèr (2023) embed a “jailbreak backdoor” into the LLM by poisoning the RLHF training data. Then, users can easily achieve jailbreak by using a trigger word like the “sudo” command in the Linux system. Wolf et al. Wolf et al. (2023) propose a theoretical approach Behavior Expectation Bounds (BEB) that investigates several limitations of alignment in LLMs, which exposes fundamental problems and emphasizes the necessity of designing reliable mechanisms to ensure AI safety.
Model Attacker: PAIR Chao et al. (2023)
No Defense SmoothLLM Robey et al. (2023) Perplexity Filter Jain et al. (2023)
GPT-3.5-turbo 76% 12% 15%
GPT-4-0125-preview 50% 25% 43%
LLaMA-2-7B-chat 4% 1% 4%
Vicuna-13B-v1.5 82% 47% 81%
Attacker: GCG Zou et al. (2023)
GPT-3.5-turbo 34% 1% 1%
GPT-4-0125-preview 1% 3% 0%
LLaMA-2-7B-chat 2% 1% 0%
Vicuna-13B-v1.5 58% 1% 1%
Attacker: JBC Albert (2023)
GPT-3.5-turbo 0% 0% 0%
GPT-4-0125-preview 0% 0% 0%
LLaMA-2-7B-chat 0% 0% 0%
Vicuna-13B-v1.5 79% 64% 79%
Table 6: An overview of defense methods under jailbreaking on listed LLMs. The evaluation method for the Attack Success Rate (ASR) indicator is consistent with JailbreakBench Chao et al. (2024). The responses of LLMs are evaluated using LLaMAGuard-7B.
5.2 Defense Methods
Only a few studies attempt to address the defense methods in the security LLMs, which can be categorized into open-source and close-source LLMs. Based on the available access to the open-source LLMs, the existing studies Deng et al. (2023); Li et al. (2024c); Robey et al. (2023) enhance the security by fine-tuning the foundation models with security instructions. For closed-source LLMs, previous works Li et al. (2023c); Wu et al. (2024); Robey et al. (2023); Jain et al. (2023) defend the risks by auditing the input prompts with various security judgment strategies. However, these simple mechanisms cannot achieve satisfactory performance, as shown in Table 6. The results demonstrate that it is difficult to completely eliminate the generation of unsafe content, regardless of the defense mechanism. And “SmoothLLM” Robey et al. (2023) performs better than “Perplexity Filter” Jain et al. (2023) but still generates over 10% unsafe content in most scenarios. It indicates that current defense methods for LLMs are in the early stages and require further exploration.
5.2.1 Defense Methods for Open-Source LLM
Robey et al. Robey et al. (2023) propose SmoothLLM, a compatible defense method that acts as a wrapper to smooth the outputs of LLMs by perturbing original attack inputs. Although effective in prohibiting the attacking intent, it also disrupts the semantics of the original input and leads to performance degradation. Deng et al. Deng et al. (2023) propose a data augment defense framework SELF-DEFENSE to defend the attacks for both unintentional and intentional scenarios. The key idea is to translate some English seed examples into a target language and then merge the language-specific corpora into the original dataset. Li et al. Li et al. (2024c) implement a fine-tuning mitigation method based on their multilingual jailbreak prompts dataset, reducing the attack success rate by 96.2%. Zhou et al. Zhou et al. (2024a) introduce Robust Prompt Optimization (RPO) to LLM security by minimizing the confrontation loss. With less impact on benign cues, the success rate of attacking GPT-4 with GUARD Jin et al. (2024) dropped from 96% to 4%.
5.2.2 Defense Methods for Closed-Source LLM
For closed-source LLM, the defense algorithms mainly focus on dealing with the model inputs. Jain et al. Jain et al. (2023) introduce a filter method, judging the danger of the input based on the perplexity. Wu et al. Wu et al. (2024) propose a simple and effective defense method SELFDEFEND, which incorporates a shadow stack to check whether harmful prompts exist in the input. Li et al. Li et al. (2023c) introduce a Rewindable Auto-regressive INference (RAIN) approach, which has the same evaluation setting as SELFDEFEND but focuses on the output of LLMs. The key component is to leverage another LLM to score the output content and determine whether the output contains harmful content, which achieves good results on LLaMA-base models.
5.3 Limitations and Future Directions on LLM Security
In previous sections, we provide a detailed introduction to the research achievements in attack and defense in terms of LLMs security. Most of the current research on LLM security is tested on those popular models (i.e., GPT-3.5, GPT-4, and LLaMA) with multilingual capabilities. Based on the summary of existing approaches for general attack and defense, we explore the following two aspects as future directions: (1) Jailbreak by targeting the multilingual ability of LLMs. (2) How to improve the robustness of LLMs in multilingualism.
5.3.1 Jailbreak by Targeting the Multilingual Ability of LLMs
Low-Resource language attack. Although LLMs already have strong multilingual capabilities, they cannot remedy the weakness of low-resource languages in the corpus. The ability of LLMs to identify harmful contents in low-resource languages becomes much worse. Existing studies Yong et al. (2023); Deng et al. (2023); Shen et al. (2024) explore a lot in low-resource languages, while only focusing on GPT-4 with incorporated translation models. Such a pipeline would inevitably introduce noise. Meanwhile, the evaluation standards should be improved to identify the security capacity of LLMs.
The asymmetry between different languages. Language security is highly culturally specific, and each language has its own security vulnerabilities, leading to asymmetry. For example, the word “uso” has insulting connotations only in Basque, thus it needs specific supervised tuning for each language, which is difficult to transfer from other languages. This contradicts the traditional multilingual methods, where knowledge between different languages can be mutually beneficial Johnson et al. (2017); Huang et al. (2023b). Xu et al. Xu et al. (2024b) empirically validate the presence of multilingual human values within LLMs, encompassing many categories of human values across multiple languages and LLM series. However, the problems of asymmetry between different languages have risen, and no existing studies about it.
Customized encryption and language rules. To bypass the security mechanism of LLMs, an innovative way to create a new language “X” by customizing some language or encryption rules Lv et al. (2024); Huang et al. (2023d); Yuan et al. (2024) from the original input. This is based on the translation paradigm, which can be viewed as a mapping function to transform character set A to character set B. Thus, exploiting the potential loopholes when aligning LLMs in different languages becomes an important problem.
Misleading information in specific languages. The prior method attempts to assign a word with a specific meaning by embedding a backdoor or directly defining it through dialogue Rando and Tramèr (2023); Jin et al. (2024). For instance, Rando et al. Li et al. (2023e) insert the word “SUDO” at the end of the query, causing the LLM to misunderstand and bypass the security defenses. It provides insight for attacking the LLMs through multilingual languages. People can introduce words or phrases with specific meanings into queries provided to LLMs. It can lead to the model misunderstanding the meaning of queries, which results in the generation of harmful information. In particular, the polysemous phenomenon that is prevalent in most languages can be used to mislead LLMs into producing prohibited content.
5.3.2 Robustness of LLMs with Multilingualism
Existing studies intend to attack LLMs to find some leaks, rather than to improve the defense ability and robustness of LLMs. However, an attack or jailbreak cannot ensure the development of LLM security. In this section, we will discuss the potential solutions for improving the security and robustness of LLMs.
Adversarial training with multilingual samples. Inspired by Mazeika et al. Mazeika et al. (2024), the proposed method enhances defense and robustness using adversarial training. Unlike fine-tuning on fixed datasets with harmful prompts, they propose a Robust Refusal Dynamic Defense to fine-tune the LLM on a dynamic pool of samples, which are continuously updated through a strong optimization-based jailbreak method. It is feasible to incorporate multilingual data into the sample pool using multilingual jailbreak methods which are more practical in attacks, making the data in the sample pool more diverse and advantageous for adversarial training. Besides, adversarial training can be introduced in the pre-training stage or fine-tuning stage of LLMs. As a result, the safety and robustness of LLMs with multilingualism can be enhanced at the root by adversarial training.
Multilingual security alignment with data augmentation. Similar to the approach of enhancing LLMs instruction-following capabilities and human values proposed by SFT and RLHF Ouyang et al. (2022), effective data augmentation is the direct way to improve multilingual alignment. Thus, it is beneficial to explore ways to construct multilingual security datasets facing the localization of security information across languages. Depending on the business or scientific requirements, the datasets can contain different harmful question pairs and languages and then fine-tune the generic LLM with customized security datasets to construct a more robust version of LLMs.
Pre-processing prompts or post-processing outputs. Another approach with fewer resources is to add proxy security efforts in small models outside of LLMs. Before constructing prompts for the LLM, we can consider maintaining a harmful vocabulary list or using another LLM as an evaluator to extract dangerous elements and judge the security of prompts, which can help detect unstable factors. The judgment outputs from LLMs with scores indicate whether the risks are identified in the original inputs and need to be rewritten, similar to RAIN Li et al. (2023c). The general goal is to force LLMs to reject harmful answers or replace them with a rule-based template.
6 Multi-Domain LLMs in Multilingual Scenarios
The remarkable capabilities of LLMs have facilitated their application across diverse domains, including finance Liu et al. (2021); Yang et al. (2023d); Liu et al. (2023e); Chen et al. (2023b); Wu et al. (2023a); Yang et al. (2023e), medicine Tian et al. (2023); Zhang et al. (2023h); Luo et al. (2024); Yang et al. (2023f); Liu et al. (2023f); Zhang et al. (2023i); Xiong et al. (2023), law Yue et al. (2023); Li et al. (2023f); Huang et al. (2023e); He et al. (2023b); Cui et al. (2023b), education Dan et al. (2023); Yu et al. (2023c), transportation Wang et al. (2024b) etc. These domain-specific LLMs have demonstrated superb capability and promising perspective within associated domains. However, these LLMs are predominantly focused on English, while fewer cater to medium or low-resource languages, which dramatically hinders the utility of LLMs for a global audience. In this chapter, we present the pioneering multilingual studies conducted in the medical and legal domainsTo the best of our knowledge, at the time we conduct our survey, there are few pertinent works in other domains such as finance and education in multilingual scenarios., and we conclude by offering a comprehensive discussion on the emerging limitations and challenges.
Model Params Base Model Training Corpus
Name Non-English proportion Translated Size Languages
BioMistral Labrak et al. (2024) 7B Mistral-7B PMC 1.25% $\times$ 3B 10
MMedLM2 Qiu et al. (2024a) 7B InterLM2-7B MMedC 58% $\times$ 25.5B 6
Apollo Wang et al. (2024a) 0.5/1.8/2/6/7B Qwen/Gemma/Yi Apollo 33.72% $\times$ 2.5B 6
Medical mT5 Garc´ıa-Ferrero et al. (2024) 0.77/3B mT5 - 66.67% $\times$ 3B 4
L2M3 Gangavarapu (2024) 70B Meditron& SeamlessM4T - 80% 0.9*5B 5
Table 7: An overview of the existing competitive LLMs with multilingualism in the medical domain. Languages denotes the number of languages included in the training corpus. Non-English proportion denotes the proportion of non-English languages in the training corpus. The training dataset of L2M3 is obtained by translating English into four low-resource languages: Telugu, Hindi, Swahili, and Arabic.
Model Benchmark
Name Translated Task Metrics Languages
BioMistralLabrak et al. (2024) - Multi-choice QA ACC en,ar,zh,fr,de,pt,ru,es
MMedLM2Qiu et al. (2024a) MMedBench Multi-choice QA Rationale eval ACC ROUGE-1/BLEU-1 en,zh,ja,fr,ru,es
ApolloWang et al. (2024a) XMedBench $\times$ Multi-choice QA ACC en,zh,fr,es,ar,hi
Medical mT5Garc´ıa-Ferrero et al. (2024) - $\times$ Seqence labeling Abstractive QA F1 score en, es, it, fr
Table 8: An overview of the existing cross-lingual medical benchmarks. For the Languages column, en, ar, zh, fr, de, pt, ru, es, ja, and hi represent English, Arabic, Chinese, French, German, Portuguese, Russian, Spanish, Japanese, Hindi, and Italian.
6.1 Medical Domain
Previous works have made substantial strides in integrating LLMs into the medical domain Tian et al. (2023); Zhang et al. (2023h); Chen et al. (2023c). In particular, Med-PaLM2 Singhal et al. (2023) notably achieves success by passing the US Medical Licensing Examination. A common practice to adapt LLMs to the medical domain is to continually train the foundation model with domain corpus as introduced in Section 3.2 Luo et al. (2022). For instance, some models such as BioGPT Luo et al. (2022), ClinicalGPT Wang et al. (2023e), PMC-LLaMA Wu et al. (2023b) and MedAlpaca Han et al. (2023) have been progressively fine-tuned foundation models on the medical related corpus, achieving dramatic medical performance enhancement over the original foundation model. In particular, ChatDoctors Yunxiang et al. (2023) resembles more of a multidisciplinary doctor, built on mixed data instead of solely medical data, capable of conducting patient-doctor dialogues, focusing on comprehensive inquiry services in the real-world scenario. In addition, in the realm of traditional Chinese medicine, HuaTuoGPT2 Chen et al. (2023d) that is fine-tuned on the Baichuan Yang et al. (2023c) with four types of data (distilled instructions/conversations from ChatGPT Ouyang et al. (2022) and real-world instructions/conversations from doctors) and capable of mimicking the diagnostic abilities of doctors provides useful medical information. The experimental results demonstrate the capabilities of HuaTuoGPT2 surpassing those of the GPT-4 Achiam et al. (2023) in the 2023 national medical licensing examination of traditional Chinese medicine. However, the aforementioned models primarily focus on a limited set of high-resource languages and show varying degrees of performance degradation when extended to other languages. Such a phenomenon renders them unreliable, incapable, and insecure for application in linguistically diverse medical environments.
6.1.1 Medical LLMs in Multilingual Scenarios
To alleviate the issue of multilingualism in the medical domain, prior studies attempt to introduce multilingual medical corpus to enhance the multilingual ability of foundation models as shown in Table 7 and Table 8, Specifically, KBioXLM Geng et al. (2023) adapts XLM-R Conneau et al. (2020b) to the medical domain, which encompasses diverse medical knowledge. However, KBioXLM explores both the training corpus and evaluation data through translation, adapted to only two languages. To further address the limited availability of data beyond English, L2M3 Gangavarapu (2024) integrates Meditron-70B with the Meta SeamlessM4T machine translation system and separately fine-tune two components on four extremely low-resource languages. Garcia et al. Garc´ıa-Ferrero et al. (2024) explore the medical LLM on encoder-decoder architecture based on mT5Xue et al. (2020). In the monolingual setting, both in-context learning (ICL) and SFT demonstrate considerable improvements compared to the foundation models Han et al. (2023); Wu et al. (2023b); Chen et al. (2023c). MMedLM2 Qiu et al. (2024a) underscores medical multilingualism by presenting a 25.5B massive cross-lingual training corpus covering a set of 6 languages. By training with this corpus, the multilingual ability of the model improves in all 6 languages, substantially surpassing the prior models and even rivaling GPT-3.5 and Gemini-1.0-pro. Meanwhile, to further assess the multilingual generalization of medical LLMs, BioMistral Labrak et al. (2024) introduces the first large-scale multilingual medical benchmark of LLMs into 7 languages. Similar to MMedLM2 Qiu et al. (2024a), the Apollo Wang et al. (2024a) introduces a training dataset “Apollo Corpora”, a benchmark “XMedBench” and a collection of models ranging from 0.5B to 7B. To investigate the effectiveness of LLMs in the multilingual medical domain, they build a benchmark “XMedBench”. The comparison on the XMedBench demonstrates that prior multilingual medical LLMs are limited by the PubMed-Central corpus which is constructed based on the translation technique. To alleviate the issue of training data, they propose a training corpus “Apollo Corpora” which is rich with high-quality medical knowledge in each language. The corpus is meticulously collected from the local language and strictly prohibits any form of translation. As a result, the Apollo-7B is fine-tuned on the “Apollo Corpora” and can achieve better performance compared with prior multilingual medical LLMs with 70B parameters.
6.1.2 Discussion and Challenges
Existing studies achieve remarkable progress in medical LLMs for the multilingual scenario, yet numerous challenges are perverse. First, given the presence of language-specific knowledge that is highly pertinent to the local cultural, historical, political, and regional backgrounds, Wang et al. Wang et al. (2024a) examine whether these language-specific medical knowledge stimulates or deteriorates with each other and investigate whether the model always outperforms its counterpart trained solely on the monolingual language corpus when further trained on the whole multilingual corpus. For instance, the ability of the Spanish model trained in 6 languages surpasses that of a model only trained on a monolingual corpus in Spanish. Despite underlying conflicts between different language-specific medical knowledge and potential biases due to varying data, the performance boost suggests that cross-lingual joint training promotes the performance of medical LLMs, shedding light on the potential of cross-lingual pre-training. Thus, further exploration into the effectiveness of the real-world and pseudo data is yet to be undertaken.
Second, the ongoing scarcity of medical data in various languages persistently hampers further advancement. Although translation can mitigate some of these issues, it may not be effective due to the complex medical terminology and challenges in precise translation. Moreover, each language might carry unique cultural and contextual differences, resulting in abundant language-specific medical knowledge intricately linked to cultural, historical, political, and regional backgrounds. For example, traditional Chinese medicine embodies a rich history deeply intertwined with its cultural heritage, constituting a distinct medical system. Enabling our systems to grasp such language-specific medical knowledge and provide tailored medical assistance for specific groups poses a formidable challenge. Hence, exploring the underlying mechanism of language-specific medical knowledge integration is a promising research direction.
Categories Details
Component Pile of Law(292 GB), Eurlex Resources2 (179 GB), Native Multi Legal Pile (112 GB), Legal MC43 (106 GB)
Text Type Distribution case law(51.4%), legal-mc4(16.6%), legislation(12.6%), contracts(9.23%), other(10.2%)
Top 5 Languages Portuguese(15.93%), German(6.29%), Spanish(6.1%), French(3.32%), Italian(2.92%)
Last 5 Languages Maltese(0.43%), Lithuanian(0.43%), Latvian(0.42%), Croatian(0.3%), Irish(0.08%)
Non-English proportion 48.27%
Total Words 86.36B
Table 9: The proportion of different languages in MultiLegalPile (689GB multilingual legal corpus).
6.2 Legal Domain
6.2.1 Legal LLMs in Multilingual Scenarios
Similar to the medical domain, the applications of LLMs in the legal domain principally center on English. Several precedent attempts such as Chatlaw Cui et al. (2023b), Lawyer LLaMA Huang et al. (2023f), SaulLM Colombo et al. (2024), and LegalBERT Chalkidis et al. (2020) expand the general ability of LLMs to the legal domain. The universal performance degradation has been observed when expanding to other languages.
To address the specific problems in the legal domain, the proposed models need to adapt the legal features, which are factual, ambiguous, structured, and timely Bansal et al. (2019); Gordon et al. (2009), compared to other domains. Brugger et al. Brugger et al. (2023) take a preliminary step in the multilingual legal Sentence Boundary Detection (SBD) covering 6 languages. Christen et al. Christen et al. (2023) conduct similar research on the multilingual Negative Scope Resolution (detecting words affected by negation cue) task and Baumgartner et al. Baumgartner et al. (2024) extend legal judgment prediction to German, French and Italian. Furthermore, for a comprehensive evaluation, Niklaus et al. Niklaus et al. (2023a) collect 11 natural language understanding legal datasets covering a total of 24 languages and 8 subdivisions, which is the first cross-lingual legal benchmark (LEXTREME) showing there is still improvement room even for the popular models like ChatGPT. After the ChatGPT series models were proposed, Nguyen et al. Nguyen et al. (2024) implement a preliminary empirical comparison between ChatGPT and GPT-4 on the “COLIEE” benchmark Rabelo et al. (2020, 2022) encompassing Japanese and English, where GPT-4 consistently surpasses its predecessor but still falls behind human performance.
For a more in-depth exploration of LLMs in multilingual legal issues, Niklaus et al. Niklaus et al. (2023b) construct a multilingual legal-domain corpus with 689GB, MultiLegalPile, whose detail is shown in Table 9. They train two PLMs based on XLM-R Conneau et al. (2020b) and the Longformer Beltagy et al. (2020) with this corpus, achieving the state-of-the-art performance on the LEXTREME Niklaus et al. (2023a) compared to XLM-R in most languages. Moreover, the English version of the model achieves state-of-the-art performance on 5 out of 7 tasks in LexGLUE Chalkidis et al. (2022b), underscoring the exceptional cross-lingual ability and legal-domain expertise. Trautmann et al. Trautmann et al. (2022) focus on employing legal prompt engineering (LPE) to enhance the capabilities of LLMs, thereby mitigating the challenges posed by the scarcity of cross-lingual legal data and the substantial computational resources required. Although improvements over baselines are observed, bridging the performance gap to match the supervised methods remains a significant endeavor.
6.2.2 Discussion and Challenges
Despite current efforts, there is still a lack of robust LLMs capable of effectively and comprehensively performing law-related multilingual tasks, highlighting the need for further exploration in this domain. Beyond the issue of data scarcity, the accumulation of language-specific legal knowledge compounds the complexity, as legal systems and jurisdictions vary significantly across regions. Given that existing LLMs already struggle with representing semantic features in low-resource languages, accurately capturing legal nuances across diverse jurisdictions poses an even greater challenge. Moreover, the temporal dimension adds complexity, as laws undergo constant revision, amendment, or abolition, necessitating that models remain continuously updated.
6.3 Limitations and Future Directions on Multi-Domain
Despite the remarkable advancements in multilingual LLMs, persistent limitations and challenges necessitate further exploration. This survey provides a brief discussion of the current limitations and potential improvement in the following two parts.
Data scarcity and translation issues. A powerful multilingual LLM, especially in specific domains, is predominantly hindered by the scarcity of domain data. Although knowledge transfer provides some relief, the issue of under-representation persists, particularly for low-resource languages, and becomes further compounded when extending to specific domains. Machine translation techniques offer a potential solution to mitigate this issue, however, they lead to new challenges Nicholas and Bhatia (2023). On one hand, machine translation systems introduce errors, particularly when handling domain-specific terminology across multiple languages. Terms or phrases that native speakers do not use may also be included in the translated corpus. On the other hand, the models suffer from comprehensively understanding and accounting for the local and cultural context of the target language, complicating the task of in-depth and high-level feature learning and capturing.
Language-Specific knowledge integration In specific contexts such as the legal or financial domain, each language harbors distinctive knowledge influenced by diverse historical, cultural, and regional backgrounds. Beyond linguistic semantics, the challenge arises in capturing these nuances among various languages and integrating language-specific domain knowledge into LLMs. For example, disparities in legal definitions between European Council and USA jurisdictions, as well as the contrast between traditional Chinese medicine and Western medicine, indicate these challenges. Current LLMs face challenges in effectively understanding such language-specific knowledge, hindering their capacity to provide customized domain-specific assistance for diverse populations. Further research is necessary to explore how LLMs can integrate and leverage this particular type of knowledge.
These limitations highlight the need for further research efforts in the following directions:
  • Development of strategies to construct high-quality, domain-specific multilingual datasets that preserve cultural context.
  • Exploration of techniques for LLMs to effectively integrate and leverage language-specific in-domain knowledge.
By addressing these two challenges, researchers can pave the way for the development of truly robust and equitable LLMs that serve a global audience in multilingual scenarios.
7 Multilingual Data Resource
Name Release Time Languages Size Sources Affiliation
Amazon intent mteb (a) 2022 49 2.02 GB Amazon Amazon
Amazon reviews Keung et al. (2020a) 2020 6 - Amazon Amazon
Bactrian-x Li et al. (2023d) 2023 51 - - MBZUAI
Biblenlp eBible.org 2024 861 581 MB Bible -
Bloom-lm Le Scao et al. (2022) 2022 364 - - SIL International
CC100 Conneau et al. (2020a) Wenzek et al. (2020) 2020 109 185 GB CC Facebook
CulturaX Nguyen et al. (2023) 2023 167 27 TB mC4, OSCAR University of Oregon
GPT-4 Prompts erfanzar 2024 5 1.05 GB GPT-4 -
Guanaco Dettmers et al. (2023) 2023 4 400 MB - University of Washington
HPLT de Gibert et al. (2024) 2023 75 22 TB CC, Internet Archive HPLT project
IWSLT 2017 Cettolo et al. (2017) 2017 10 4.24 GB TED Talks FBK
mC4 Raffel et al. (2019) 2019 101 9.7 TB CC Google
Mewsli-x Ruder et al. (2021a) 2021 50 285 MB WikiNews and Wikipedia DeepMind
Minds14 Gerz et al. (2021) 2021 12 471 MB Banking Assistant PolyAI Limited
Miracl Zhang et al. (2022b) 2023 18 - Wikipedia University of Waterloo
MLDR Chen et al. (2024b) 2024 13 - Wikipeida, Wudao, mC4 BAAI
MMedC Qiu et al. (2024b) 2024 6 105 GB Multiple SJTU
MQA De Bruyn et al. (2021) 2021 36 122 GB CC CLiPS Research
Multi-sentiments tyqiangz 2022 12 141 MB Multiple Sources -
Multiconer Fetahu et al. (2023a) Fetahu et al. (2023b) 2023 12 338 MB Wikipedia Amazon
Nomiracl Thakur et al. (2023) Thakur et al. (2024) 2023 16 - Wikipedia University of Waterloo
Open Subtitles Lison and Tiedemann (2016) 2023 58 273 MB Subtitles -
OSCAR Abadji et al. (2022) 2020 166 6.3 TB CC University of Orego
Para-pat Soares et al. (2020) 2020 15 2.57 GB Google Cloud University of Sheffield
Project Gutenberg Gutenberg 2023 11 14.4 GB eBook Project Gutenberg
ShareGPT RyokoAI 2023 5 1.08 GB GPT / Human Ronso
SREDFM Huguet Cabot et al. (2023) 2023 18 8.29 GB Wikidata, Wikipedia Babelscape
TED Talks Qi et al. (2018) 2018 55 - TED Talks CMU
TED-talks-iwslt Cettolo et al. (2012) 2012 104 25 GB TED Talks FBK
Toxi-text FredZhang 2023 55 1.96 GB - -
UD Nivre et al. (2020) 2023 102 2.19 GB - Universal Dependencies
Wikiann Pan et al. (2017) Rahimi et al. (2019) 2019 173 143 MB Wikipedia RPI
Wikipedia Foundation 2024 322 71.8 GB Wikipedia Wikimedia Foundation
Wit Base Srinivasan et al. (2021) 2021 105 5.15 GB Wikipedia Google
xP3 Muennighoff et al. (2022a) 2022 277 - - Cohere For AI
Table 10: An overview and statistic detail of the representative multilingual data resource. We only include the large-size datasets with much more supported languages.
LLMs are data-driven, thus the impressive learning capabilities of LLMs stem from their massive model sizes and extensive training datasets, which have been proven in high-resource languages. However, English stands as the closest approximation to the lingua franca, wielding dominance across various domains. With the largest number of total speakers, its prominence extends far and wide Pereltsvaig (2020), where English reigns as the primary language of the internet Schneider (2018); Pennycook (2002). Meanwhile, English is the main language used in the higher economic status countries of the world, such as America, Britain, and other Western countries Joshi et al. (2020). Therefore, existing data resources focus on the English-centric, which came at the expense of regional and indigenous languages, exacerbating language endangerment and economic marginalization Dodge et al. (2021). Due to the lack of resources, this situation deeply restricts the development of multilingual models, and it is a vicious circle Caswell et al. (2020). Moreover, low-resource languages suffer from lower quality, due to mislabeling or inadequate representation of native usage. This situation is especially prevalent with web-crawled data which predominantly consists of pornographic, nonsensical, or non-linguistic content Kreutzer et al. (2022).
As shown in Table 10, we collect as much large-scale multilingual data resource as is reliable. We can observe that the scale of existing multilingual resources is much smaller than that of English monolingual resources (only four datasets with the TB level). And the low-/medium-resource languages typically derive data from a narrower range of sources compared to their high-resource counterparts Nekoto et al. (2020). The available data mainly originates from sources, such as Wikipedia, the Bible, and Common Crawl. In addition, these data suffer from bias and fairness, which we will present in Section 9. For instance, a gender bias exists in Wikipedia, with studies revealing a persistently low percentage of women editing articles Callahan and Herring (2011). The vicious circle of multilingual data includes the issue of open source issues. Due to the high cost of high-quality multilingual resources, researchers and companies are reluctant to share resources in the open-resource community (i.e., “data island”) Peris et al. (2023); Yao et al. (2024); Yan et al. (2024); Kandpal et al. (2022). This situation results in multilingual research staying at the data level and ignoring the competition in the model paradigm. The united governments, companies, and researchers must start a virtuous cycle of multilingual data resources. Access to abundant, meticulously collected datasets in a language empowers researchers and developers to construct models and benchmarks. The abundance of models and benchmarks, in turn, fosters increased publication, facilitates communication, and promotes real-world application scenarios for companies. These outputs have the potential to attract more users, while government-mandated guidelines help generate non-toxic data, which can be repurposed for further research and development.
8 Multilingual Benchmark and Evaluation
Name Release Time Languages Parallel Type Affiliation
Afrisent Muhammad et al. (2023) 2023 14 $\times$ Sentiment Analysis U.Porto
ASPEN Razumovskaia et al. (2024) 2022 31 Story Planning Cambridge University
BELEBELE Bandarkar et al. (2023) 2023 122 QA Meta
BioInstructQA BioMistral 2024 7 Question Answering Avignon Université
Bucc-bitext-mining mteb (b) 2022 5 $\times$ Mixed HuggingFace
Cross-Sum Bhattacharjee et al. (2023) 2021 45 Summarization BUET
Crossmodal-3600 Thapliyal et al. (2022) 2022 36 Image Captioning Google
Exams Hardalov et al. (2020) 2020 15 Examination Sofia University
Fairlex Chalkidis et al. (2022a) 2022 5 $\times$ Fairness University of Copenhagen
FLORES-200 Team (2022) Goyal et al. (2022) Guzmán et al. (2019) 2022 200 Machine Translation Meta
GEOMLAMA Yin et al. (2022) 2022 5 Knowledge Diversity University of California
Humaneval-XL Peng et al. (2024) 2024 23 Code Generation University of Copenhagen
M-Hellaswag Dac Lai et al. (2023) 2023 35 Commonsense NLI University of Oregon
M-MMLU Dac Lai et al. (2023) 2023 35 NLU University of Oregon
M3Exam Zhang et al. (2023d) 2023 9 $\times$ Examination DAMO Academy
M3LS Verma et al. (2023) 2023 20 $\times$ Summarization IITs
MARC Keung et al. (2020b) 2020 6 $\times$ Text Classification Amazon
MasakhaNER Adelani et al. (2021) 2021 10 Named Entity Recognition Masakhane
Masakhanews Adelani et al. (2023a) 2023 16 $\times$ News Topic Classification Masakhane
MASSIVE FitzGerald et al. (2022) 2022 51 Intent Recognition, Slot Filling Amazon
MaXM Changpinyo et al. (2023) 2023 7 Visual Question Answering Google
MEGA Ahuja et al. (2023) 2023 70 $\times$ Mixed UW
MEGAVerse Ahuja et al. (2024) 2023 83 $\times$ Mixed Microsoft
Mela Zhang et al. (2024b) 2022 10 $\times$ Linguistic Acceptability SJTU
MGSM Shi et al. (2022a) 2022 11 Mathematical Reasoning Google
MLQA Lewis et al. (2020b) 2020 7 Question Answering Facebook
MMedBench Qiu et al. (2024c) 2023 6 $\times$ Medical SJTU
Multi-CoNER Malmasi et al. (2022) 2022 11 $\times$ Named Entity Recognition Amazon
MULTIEURLEXDOC Chalkidis et al. (2021) 2021 23 Legal Topic Classification University of Copenhagen
Multilingual-Fig-QA Kabra et al. (2023) 2023 7 $\times$ NLI CMU
NusaX Winata et al. (2023b) 2023 10 Mixed Bloomberg
ODEX Wang et al. (2023d) 2022 14 Code Generation CMU
OPUS-100 Zhang et al. (2020) 2020 100 Machine Translation University of Edinburgh
Paws-X Yang et al. (2019) 2019 7 Paraphrase Identification Google
Pmindiasum Urlana et al. (2023) 2023 14 Summarization IIIT Hyderabad
PRESTO Goel et al. (2023) 2023 6 $\times$ Conversational Parsing University of Rochester
SEAHORSE Clark et al. (2023) 2023 6 $\times$ Summarization DeepMind
Sib200 Adelani et al. (2023b) 2024 205 Topic Classification UCL
SMiLER Seganti et al. (2021) 2021 14 $\times$ Entity and Relation Extraction Samsung
STSB-multi-mt May (2021) 2021 10 Text Similarity -
Tatoeba-mt Tiedemann (2020) 2020 93 Machine Translation Tatoeba.org
Tydip Srinivasan and Choi (2022) 2022 9 $\times$ Identify Politeness Levels UT-Austin
TyDiQA Clark et al. (2020) 2020 11 $\times$ Question Answering Google
Universal Dependencies Nivre et al. (2020) 2022 146 $\times$ Parser -
X-CLAIM Mittal et al. (2023) 2023 6 $\times$ Realworld Claims MBZUAI
X-RiSAWOZ Moradshahi et al. (2023) 2023 6 Dialogue Utterances Stanford University
XCOPA Ponti et al. (2020) 2020 12 Commonsense Reasoning Cambridge University
XCSQA Lin et al. (2021b) 2021 11 Commonsense Reasoning USC
xDial-Eval Zhang et al. (2023e) 2023 10 Dialogues NUS
XGLUE Liang et al. (2020) 2020 - $\times$ Mixed Microsoft
XL-SUM Hasan et al. (2021) 2021 44 $\times$ Summarization BUET
XNLI Conneau et al. (2018) 2018 15 NLI NYU
XQuAD Artetxe et al. (2020) 2020 10 Question Answering University of the Basque Country
XSEMPLR Zhang et al. (2023f) 2023 22 $\times$ Semantic Parsing PSU
XStoryCloze Lin et al. (2022) 2022 11 Commonsense Reasoning Meta
XTREME Hu et al. (2020) 2020 40 Mixed CMU
XTREME-R Ruder et al. (2021b) 2021 50 Mixed DeepMind
XWinograd Tikhonov and Ryabinin (2021) 2021 6 $\times$ Commonsense Reasoning Yandex
XCSR Lin et al. (2021c) 2021 16 Commonsense Reasoning USC
Table 11: An overview of multilingual benchmarks with more than four supported languages after 2018 in which mBERT Devlin et al. (2018) was proposed. NLU and NLI denote the natural language understanding and the natural language inference task, respectively.
With the emergence of novel models and algorithms, researchers inevitably scrutinize the capabilities by evaluating their performance on specific and challenging tasks Chang et al. (2024). Recently, LLMs have garnered significant interest in the academia and industry. The remarkable performance of LLMs has been proven on a variety of tasks, showing strong universality compared with prior PLMs limited to specific tasks. As shown in Table 11, to the best of our knowledge, we list the representative multilingual benchmarks after 2018 in which mBERT Devlin et al. (2018) was proposed. From the statistical results, we can observe many kinds of existing benchmarks, but there remain some issues with these benchmarks described as follows.
  • Lack of task types. Most of the multilingual benchmarks focus on understanding tasks and lack generation tasks that mainly consist of machine translation and summarization. However, LLMs are generative paradigms, and generation tasks closer to the real world should be introduced to evaluate the effectiveness of LLMs in multilingual scenarios. In addition, the evaluation dimensions of LLMs need to be richer today, and there is a lack of evaluation tasks for LLMs, such as safety, agents, social simulation, etc.
  • Language culture and domain. Existing multilingual benchmarks often rely on machine-translated text, which may contain errors or terms not commonly used by native language speakers. The benchmark of native usage with language culture habits is urgently required. Besides, the challenges within different language environments vary significantly, necessitating a thorough exploration of multilingual and multi-domain issues.
  • Unified framework. The number of benchmarks is sufficient but there is a lack of an authoritative and unified evaluation framework, which can be updated over time, and the evaluation aspect is more comprehensive based on the proposed framework. This phenomenon can be attributed to the dominance of English and individuals’ focus on their native languages. To address this issue, collaborative efforts from the multilingual community are essential.
  • Data leakage. The main differences between LLMs lie in the training data and model size. The existing systems primarily prioritize real focus on user experience (open test), rather than assessing effectiveness through closed test sets that align with fair training data. Thus, there is a potential for data leakage as models may have inadvertently learned from the test set, especially the closed LLMs. It requires the evaluation methods to adopt a more secure strategy to reduce the risk of data leakage.
  • Evaluation methods. There are limitations in existing evaluation methods, particularly in generating tasks such as BLEU and ROUGE, thus manual evaluation is more reliable than automatic evaluation. However, due to the diversity of multilingual tasks, performing manual evaluation requires numerous language experts, which leads to increased costs and makes it challenging to achieve multilingual tasks. Therefore, reliable automatic evaluation methods are needed, which would also advance the development of evaluation techniques.
9 Bias and Fairness
9.1 Bias Categories
The bias of LLMs in the multilingual scenario can be divided into two categories: language bias and demographic bias Xu et al. (2024a). Intuitively, the former is due to the imbalance of available training corpus for different languages Wei et al. (2023a); Nicholas and Bhatia (2023), where English possesses the most text corpus Dodge et al. (2021); Penedo et al. (2024); Gao et al. (2020); Soldaini et al. (2024), resulting in the performance degradation of the LLMs when generalized to other language settings Zhang et al. (2023j, 2024c). The demographic bias occurs due to embedded biases and misinformation on the internet, leading to LLMs unavoidably inheriting demographic biases across gender, race, and political backgrounds Ferrara (2023); Wang et al. (2022c); Levy et al. (2023); Yeh et al. (2023). This exacerbates existing inequalities, perpetuates stereotypes, and reinforces discrimination.
The core of multilingual LLM research is to improve the language modeling ability in other medium/low-resource languages while maintaining competence in English. Concerning demographic biases, previous attempts to mitigate bias and align LLMs with human values have primarily concentrated on English Hutchinson et al. (2020); Nadeem et al. (2021); Wan et al. (2023). Consequently, bias and ethical issues persist in other languages, potentially leading to significant negative impacts for non-English-speaking users.
Name Languages Type of Bias Debias Object Metrics
RTP-LX de Wynter et al. (2024) 28 Bias, Identity attack, Insult Microaggression, Self-harm Sexual content, Toxicity, Violence Small LM Large LM Exact Label Match Interrater Reliability de Wynter et al. (2024)
MGB Yu et al. (2024) zh,en,de,pt,es Gender Bias Masked LM MBE, MGL, LSG, MSG Yu et al. (2024)
MIBs Zhao et al. (2020) en,es,de,fr Gender Bias, Occupation Bias Word Embedding inBias Zhao et al. (2020)
MozArt Piqueras and Søgaard (2022) en,de,es,fr Gender Bias, Language Masked LM Close Test
Table 12: An overview of the multilingual bias evaluation datasets. The 28 languages supported by RTP-LX are Arabic, Hebrew, Indonesian, Danish, Norwegian, Swedish, Dutch, English, German, Russian, Ukrainian, Czech, Polish, Serbian, Bosnian, Croatian, Montenegrin, Spanish, Portuguese, French, Italian, Hindi, Thai, Kiswahili, Chinese, Japanese, Korean, Turkish, Finnish, and Hungarian.
9.2 Multilingual Debias
Language bias of LLMs persists in the multilingual scenario as a consequence of the dominance of English resources and the insufficiency of other languages on the internet. To enhance the model ability on low-resource languages, a common practice is to incorporate large-scale data Gao et al. (2020); Dodge et al. (2021) for training. The extensive training data facilitates language transfer, especially among typologically similar languages. Furthermore, the strategies such as curriculum learning Wei et al. (2023a) and up-sampling Lin et al. (2022); Shliazhko et al. (2022); Labrak et al. (2024) progressively increase the proportion of non-English resource. These techniques expose LLMs to a wider range of languages while maximizing the utilization of existing data.
To mitigate demographic bias in the multilingual scenario, Zhao et al. Zhao et al. (2020) extend word embedding bias to the cross-lingual and Piqueras et al. Piqueras and Søgaard (2022) evaluate group bias of three pre-trained LM (mBERT, XLM-R, and mT5) on four languages (En, Es, De, and Fr). Besides, Vashishtha et al. Vashishtha et al. (2023) extend debiasing strategies such as counterfactual data augmentation and self-debias to non-English languages, revealing a greater potential for debiasing and generalization among linguistically similar languages. However, they only investigate a few Indian languages, without comprehensive mitigation strategies for broader language groups. For a more comprehensive evaluation, De et al. de Wynter et al. (2024) introduce RTP-LX, a dataset designed for identifying culture-specific toxic languages with much wider coverage (28 languages and 8 different classes). Experimental results on up-to-date LLMs (Mistral Jiang et al. (2023), Gemma Team et al. (2024), GPT-4 Yuan et al. (2024), etc) demonstrate that even the popular models still struggle to judge history or content-dependent toxic content. Moreover, Lin et al. Lin et al. (2022) and Shliazhko et al. Shliazhko et al. (2022) analyze safety and bias in multilingual PLMs (XGLM and mGPT), respectively. They observe that multilingual PLMs pronounce gender bias in certain occupations, while few-shot learning has minimal impact on performance improvement. To analyze gender bias robustly, Yu et al. Yu et al. (2024) propose a novel model-based approach to generate sentence pairs. Based on mBERT Devlin et al. (2018), Reusens et al. Reusens et al. (2023) investigate cross-lingual transferability of debias techniques on 4 languages and stimulate cross-lingual debiasing effectiveness with additional pre-training.
9.3 Limitations and Discussion
To mitigate the cross-lingual bias is a promising research question and it has profound implications for the welfare, fairness, and esteem of numerous social and racial groups. Due to the scarcity of high-quality data in low-resource languages and the absence of pertinent evaluation benchmarks, effective bias detection and elimination remain largely unexplored, underscoring the necessity and imperative for future research. Besides, machine translation can mitigate the issue of low-resource languages via pseudo data generation Prabhumoye et al. (2018); Hoang et al. (2018) but it omits the cultural context or fails to capture the cultural nuances of a specific language Brislin (1970). The local and cultural backgrounds are critical to prejudice and hate speech. Thus, leveraging raw corpus in the local original context is important for detecting toxic content by native speakers.
10 Conclusion and Future Directions
In this paper, we summarize the existing representative research efforts on LLMs in the multilingual scenario from multiple perspectives. We first rethink the transitions between previous and current research on pre-trained language models. Based on the main aspect, the survey is divided into several sections from the view of training paradigms, inference strategies, security, multi-domain, data resources, and benchmark evaluation. Besides, we appeal to the research community for bias and fairness when exploring multilingual models. We also discuss several urgent challenges related to each investigated aspect and provide reflections and potential solutions for future work. Finally, considering the rapid growth of LLM research, we establish a continuously updated repository to provide relevant literature with the latest advancements of LLMs in the multilingual scenario.
In conclusion, LLMs have greatly contributed to the advancement of multilingual applications, progressing toward the goal of user-oriented. However, the existing technologies and algorithms in various multilingual tasks still fall short of expectations, which makes it difficult to meet practical standards. Aiming for language-fair AI, extensive research efforts are required to adapt LLMs for multilingual tasks much more feasible. We summarize the suggestions for both academic and industry as they build, study, and regulate LLMs as follows:
  • Sustainable language adaptation. The limited data resources of various languages restrict the number of supported languages with the initial model pre-training. An ideal situation is to use the newly available language data to improve the performance and supported languages of LLMs. Although mammalian brains could protect previously acquired knowledge through cortical loops to avoid catastrophic forgetting, the neural network models lack this capability. Therefore, sustainably achieving language adaptation for LLMs is not trivial. The longstanding goal of LLMs with multilingualism is to achieve good performance among multiple languages for all tasks in an incremental learning paradigm.
  • Universal multilingual paradigm. The existing studies mainly focus on leveraging parameter-tuning techniques and prompt engineering to explore the potential multilingual capabilities of LLMs. Aiming for a universal multilingual paradigm based on LLMs, it is beneficial to investigate the potential mechanisms without additional training to effectively address language-specific issues, such as code-switching, multilingual jailbreaking, cross-domain adaptation, etc.
  • Comprehensive and authoritative evaluation. The majority popular mainly focus on English and their native languages because of regional and linguistic restrictions, which poses a challenge in bridging the language barrier. To mitigate the language barrier issue, an urgent requirement for the multilingual community is to construct a comprehensive and authoritative benchmark to evaluate the multilingual capabilities of LLMs with various aspects, including language culture, multilingual security, multilingual reasoning, domain knowledge in native languages, etc. This can be achieved by a reasonable combination of multiple benchmarks or guidelines initiated by linguistic experts from corresponding language communities.
  • Bias impact with multilingualism. Existing LLMs inherit biases from the training corpus because of a lack of feasible data management/processing, which poses generation risks. Besides, the high proportion of Western languages in the training corpus exacerbates the bias issues from a culturally insensitive generation aspect. How to enable LLMs to avoid generating biased/risky content and to possess the ability to generate cultural concepts within different languages are important and meaningful to achieve language-fair technology.
Acknowledgment
We sincerely thank the valuable feedback from every domain expert, including Junpeng Liu (Dalian University of Technology, China), Jinsong Su (Xiamen University, China), Deyi Xiong (Tianjin University, China), and the anonymous others. In particular, Chaoqun Liu (Nanyang Technological University, Singapore) provides valuable thoughts and contributes part of the implementation of the multilingual inference strategies.